### v1

# LoRA-Mixer: Coordinate Modular LoRA Experts Through Serial Attention Routing

Abstract
Recent efforts to combine low-rank adaptation (LoRA) with mixture-of-experts (MoE) for adapting large language models (LLMs) to multiple tasks, yet exhibit prevailing limitations: they either swap entire attention/feed-forward layers for switch experts or bolt on parallel expert branches, diluting parameter efficiency and task fidelity. We propose the LoRA-Mixer, a modular and lightweight MoE framework that integrates LoRA experts. Our core innovation lies in replacing the projection matrices of the attention module’s input/output linear layers with dynamically routed, task-specific LoRA experts. This design ensures seamless compatibility with diverse foundation models—including transformers and state space models (SSMs)—by leveraging their inherent linear projection structures. The framework supports two operational paradigms: (1) joint optimization of LoRA experts and routing mechanisms via a novel hard-soft routing strategy, or (2) direct deployment of pre-trained, frozen LoRA modules sourced from external repositories. To enable robust router training with limited data while ensuring stable routing decisions and maximizing expert reuse, we introduce an adaptive Specialization Balance Loss (SBL) that jointly optimizes expert balance and task-specific alignment. Extensive experiments on seven benchmark datasets, including MedQA, CoLA, SST-2, GSM8K, ARC-E, ARC-C, and HumanEval, demonstrate the effectiveness of LoRA-Mixer. On datasets such as GSM8K, HumanEval, and MedQA, LoRA-Mixer achieves significant improvements of 7.61%, 4.88%, and 3.08% over the base models, respectively. Compared with the state-of-the-art methods, LoRA-Mixer achieves additional improvements of 1.09%, 1.45%, and 1.68%, respectively, using only 48% of the parameters, demonstrating its efficiency and strong performance.
1 Introduction
Large Language Models (LLMs) have achieved unprecedented proficiency in general-purpose reasoning and generation, yet their adaptation to specialized downstream domains remains computationally prohibitive, requiring significant resources for full-scale fine-tuning [1, 2]. To mitigate these resource demands, parameter-efficient fine-tuning (PEFT) methods [37, 38, 39, 40, 41, 42] have emerged as a scalable paradigm for task-specific adaptation. Among these, Low-Rank Adaptation (LoRA) [4, 5] has demonstrated particular efficacy, operating through low-rank decomposition of updates to the pretrained weights—enabling efficient tuning with minimal parameter overhead. Recent work has explored modularly composing independently trained LoRA modules as a promising strategy for multitask adaptation; however, naive composition can result in interference between task-specific subspaces, limiting their synergistic potential[15, 14]. This limitation has motivated exploration of mixture-of-experts (MoE) architecture [3, 6], which treat each task-specific LoRA as an expert and sparsely activate and fuse the experts. Recent studies demonstrate promising directions in hybrid Lo-
Figure 1: MoE and LoRA-MoE integration methods. (1) and (2): replace the attention or feedforward layers with switch experts; (3) and (4): introduce LoRA experts branches in parallel with the attention or feedforward layers, and fuse the output into the main branch. Our LoRA-Mixer (right) applies mixture of LoRA experts to the projection layers, which can effectively leverages the attention mechanism.
RA-MoE frameworks [14, 13, 15, 16, 35, 34, 23, 24], aiming to enhance model performance on complex tasks across multi-domain datasets while preserving the parameter efficiency of fine-tuning.
The central challenge in composing multiple pre-trained LoRAs lies in effectively synergizing them—achieving enhanced performance across constituent tasks while minimizing training overhead and preserving their distinct, task-specific characteristics. Conventional methods integrating LoRA into mixture-of-experts (MoE) architectures typically follow one of two paradigms: (1) directly substituting attention or feedforward layers with LoRA-based switchable experts [13, 23], emulating the classical MoE structure [6, 3]; or (2) introducing parallel branches of LoRA experts whose outputs are subsequently fused into the primary model pipeline [14, 15], as illustrated in Fig.1. While these strategies have demonstrated promising empirical results [14, 35, 52, 53, 54, 55], they still encounter fundamental limitations. Specifically, the vanilla MoE-inspired paradigm necessitates joint training across all expert modules, significantly increasing training data demands and restricting the modular reuse and transferability of pre-trained LoRAs. Conversely, the parallel LoRA-expert approach circumvents the inherent attention or state transition mechanisms, resulting in simplistic output fusion and suboptimal overall integration. Additionally, auxiliary losses commonly employed for routing optimization inherently promote uniform load balancing among experts, diminishing the capacity for nuanced task-awareness [13]. These limitations motivate the development of a more flexible, plug-and-play framework that is model-architecture agnostic, compatible with both Transformer and state-space model (SSM) architectures [7], capable of training an efficient and discriminative routing mechanism with minimal computational and data demands, and maximizing the reuse and transferability of independently pre-trained LoRA modules.
In this paper, we introduce LoRA-Mixer, a novel framework designed to efficiently synergize multiple pre-trained LoRA modules by treating them as dynamic, pluggable memory cells. LoRA-Mixer equips the linear projection layers of the original model with mixed LoRA experts, enabling these experts to directly leverage the effectiveness of the core attention or state-transition mechanisms. LoRA-Mixer supports LoRA modules sourced from external repositories, independently trained, or jointly trained through hard routing strategies, allowing seamless plug-and-play usage across various tasks and domains. Importantly, our method significantly reduces the necessity for training data or extensive re-adaptation, requiring only minimal additional data to effectively train the routing mechanism. Consequently, LoRA-Mixer is particularly suitable for constructing large-scale, modular language models characterized by task-specific memory, computational efficiency, and strong transferability. To further enhance efficiency and maintain routing effectiveness, we propose a novel Router Specialization Balancing Loss (RSBL). RSL aligns routing decisions with token-level expert usage, maintaining moderate entropy to encourage exploratory behavior. During inference, we employ sparse top-$K$ fusion, effectively balancing computational cost and scalability without compromising expert selectivity. Extensive evaluations conducted on seven benchmark datasets—MedQA, CoLA, SST-2, GSM8K, ARC-E, ARC-C, and HumanEval—demonstrate that integrating LoRA-Mixer substantially improves model performance across all evaluated tasks. Additionally, cross-domain experiments confirm the versatility and adaptability of our proposed framework. Compared to state-
Figure 2: The overall architecture of LoRA-Mixer. LoRA-Mixer is applied to the linear projection layers in serial with the Attention and SSM modules and support all major LLM structures. LoRA-Mixer reueses the LoRA experts sourced from Internet, trained individually or jointly trained using hard routing. The routing training is guided by RSL loss for balancing experts loads and specificity.
of-the-art approaches, LoRA-Mixer achieves significant efficiency, using only 48% of the training parameters while improving performance by 1.09%, 1.45%, 1.19%, and 0.85% on GSM8K, HumanEval, SST-2, and CoLA benchmarks, respectively.
2 Related Work
PEFT For LLMs Low-rank adaptation (LoRA) [12, 11, 4, 10, 47, 49, 50, 51] effectively fine-tunes large models by learning a low-rank matrix and freezing the original weights. While effective for a single task, its task-specific nature limits generalization. Recent work combines LoRA with mixture of experts (MoE) [13, 14, 25, 15, 16] to achieve dynamic adaptation. For example, MixLoRA [13] uses LoRA experts for top-k routing in FFN, improving multi-task performance but suffering from gradient entanglement issues. MoLE [14] combines LoRA layers via gating but lacks sparse routing. LoraHub [15] performs gradient-free few-shot combination of LoRA modules for unseen tasks but struggles with complex semantics due to lack of gradient optimization and dynamic routing. Other methods [18, 19, 45, 46, 48, 60] explore flexible routing mechanisms to improve the model’s adaptability. However, these methods usually introduce additional routing networks or optimization targets, resulting in instability during training, limiting their application in actual multi-task or low-resource scenarios.
Mixture of Experts In recent years, the mixture of experts (MoE) architecture has attracted much attention as a promising LLM expansion paradigm. By selectively activating a subset of expert modules for each input, MoE allows the model to scale capacity without linearly increasing the amount of computation. As a result, more and more large models have adopted MoE, including GLaM[20], Switch Transformers[6], and the recent DeepSeek series[21, 22]. These advances indicate that MoE is becoming a mainstream architectural trend in the development of next-generation base models. Among them, GLaM[20] and Switch Transformer[6] build a mixture of experts (MoE) model in the FFN module, and use a sparse activation mixed expert architecture to expand the model capacity and achieve better performance. LLaVA-MoLE[23] uses a top-1 strategy to route tokens to domain-specific expert models, thereby alleviating data conflicts and achieving continuous performance improvements over the ordinary LoRA baseline. LoRAMoE[24] uses routers to integrate LoRA experts while retaining general knowledge. HMoRA[35] combines the layered fine-tuning methods of MoE and LoRA, and gradually switches the routing strategy as the number of layers increases. MoLA[34] assigns different numbers of experts at different levels, proving that deeper layers often require more experts. In our method, we use mixed LoRA experts in the projection layer and optimize the load loss. LoRA-Mixer is a more fine-grained MoE construction method that is independent of the architecture and can simultaneously maintain expert expertise, computational sparsity, and routing adaptability. In addition, LoRA-Mixer can reuse existing LoRA modules and only requires very little data to train routing. While saving computing resources, it achieves the expansion of model capacity and the improvement of generalization ability.
3 Method
In this section, we introduce LoRA-Mixer, a flexible and pluggable mixture of experts (MoE) framework for combing multiple LoRA experts of LLMs.
3.1 Preliminaries
LoRA [4] is a parameter-efficient fine-tuning method that adapts large pre-trained models by introducing low-rank updates to the original weight matrices, instead of updating them directly. Given a pre-trained weight matrix $W_{0}\in\mathbb{R}^{d_{\text{out}}\times d_{\text{in}}}$, LoRA freezes $W_{0}$ and introduces a trainable low-rank decomposition $\Delta W=BA$, where $A\in\mathbb{R}^{r\times d_{\text{in}}}\,,\ B\in\mathbb{R}^{d_{\text{out}}\times r}\,,\ r\ll\min(d_{\text{in}},d_{\text{out}})$. The adapted transformation becomes:
$y=(W_{0}+BA)\mathbf{x}=W_{0}x+B(A\mathbf{x}),$ (1)
where $\mathbf{x}\in\mathbb{R}^{d_{\text{in}}}$ is the input token representation. This technique significantly reduces trainable parameters to $r(d_{\text{in}}+d_{\text{out}})$, enabling scalable fine-tuning with limited resources.
Mixture-of-Experts [3] is a sparse neural architecture where each input token is processed by a small subset of expert networks. Given $K$ experts and a router that produces a score vector $G(\mathbf{x})\in\mathbb{R}^{K}$, a softmax is applied to obtain the routing distribution:
$p_{i}(\mathbf{x})=\frac{\exp(G_{i}(\mathbf{x}))}{\sum_{j=1}^{K}\exp(G_{j}(\mathbf{x}))},\quad i=1,\dots,K.$ (2)
The top-$k$ experts are selected based on $p_{i}(x)$, and the final output is computed as a weighted sum over their outputs:
$\text{MoE}(\mathbf{x})=\sum_{i=1}^{K}\mathbb{I}[i\in\text{TopK}(p(\mathbf{x}))]\cdot p_{i}(\mathbf{x})\cdot\text{Expert}_{i}(\mathbf{x})$ (3)
This design allows MoE to reduce computational cost while enabling experts to specialize on different input patterns.
3.2 LoRA-Mixer for Compositing LoRAs
Combining independently trained LoRA modules for multi-task adaptation provides a promising approach to provide LLMs with cross-domain composition capabilities. For example, we can fuse mathematics- and medicine-specific LoRA to enable LLMs to have both stronger mathematical reasoning capabilities and medical-specific knowledge to solve complex cross-domain queries.
Our proposed LoRA-Mixer implements this mechanism by treating each pre-trained LoRA module as an expert and learning a routing function $\mathcal{F}_{\text{route}}$ that dynamically fuses these experts based on the input semantics. The routing mechanism is lightweight and data-efficient, and can achieve task awareness with only a small amount of additional training. LoRA-Mixer uses a set of $E$ low-rank experts and a router $\alpha(x)\in\mathbb{R}^{E}$ to enhance the pre-trained projection matrix $W\in\mathbb{R}^{d_{\text{out}}\times d_{\text{in}}}$. Each expert is parameterized as $\Delta W^{(e)}=A^{(e)}B^{(e)}$, where $A^{(e)}\in\mathbb{R}^{d_{\text{out}}\times r}$ and $B^{(e)}\in\mathbb{R}^{r\times d_{\text{in}}}$. The output of LoRA-Mixer is:
$\mathbf{y}=W\mathbf{x}+\mathcal{F}_{\text{route}}\left(\left\{\alpha_{e}(\mathbf{x})\cdot\Delta W^{(e)}\mathbf{x}\right\}_{e=1}^{E}\right)$ (4)
where, $\mathcal{F}_{\text{route}}(\cdot)$ represents the routing function output by the fusion expert. The output will be passed to the subsequent attention module or state-space module, enabling it to directly influence the core representation learning path. This strategy ensures that LoRA-Mixer acts at the most expressive point of the model - the projection layer - without disrupting the underlying architecture.
LoRA Experts Acquirement. Our proposed LoRA-Mixer framework is highly flexible and supports the integration of LoRA modules from diverse sources. In one common scenario, users may download pre-trained LoRA adapters from public repositories such as LoRAHub [15], which currently hosts 196 high-quality LoRA modules across a wide range of domains. These can be directly composed using LoRA-Mixer with minimal additional data. Alternatively, users may independently train domain-specific LoRA modules tailored to their own datasets. For scenarios requiring joint training of multiple LoRA modules on a heterogeneous, labeled dataset, LoRA-Mixer further supports a hard-routing strategy. Specifically, we fix the routing module and apply a deterministic routing scheme based on known domain labels. Given a domain ID $d\in\{1,\dots,K\}$ associated with each training instance, all tokens within the sample are routed exclusively to expert $d$. This design enables efficient joint optimization while maintaining expert modularity. Collectively, these capabilities make LoRA-Mixer a versatile and scalable framework for composing heterogeneous LoRA modules. The overall architecture is illustrated in Figure 2.
3.3 Specialization Balance Loss for Routing Optimization
The next step is to optimize the expert router. Although previous studies  [22, 21, 13] introduced auxiliary losses to align the average gating score with the expert utilization to promote load balancing, we observed that this approach overemphasized consistency, resulting in an overly balanced distribution of experts. In this case, all experts are forced to be used equally regardless of the input semantics. This hinders effective routing and often requires more training data. To ensure that the expert load is balanced and the routing is input-aware, we propose an improved optimization objective called Route-Specialization Balance Loss (RSL).
We introduce a selectivity-aware regularization term that regulates the entropy of routing distribution to enhance the auxiliary loss, guiding routers to make more discriminative expert choices instead of blindly averaging activations. Formally, let $\bar{p}_{i}$ denote the average soft route score (across tokens) and $\bar{f}_{i}$ denote the normalized score of the token assigned to expert $i$ in the first $k$ routes. The RSL loss function is defined as:
$\mathcal{L}_{\text{RSL}}=\alpha\cdot\sum_{i=1}^{K}\bar{p}_{i}\cdot\bar{f}_{i}-\lambda\cdot\mathbb{E}_{x\sim\mathcal{D}}\left[\mathcal{H}(p(\mathbf{x}))\right],$ (5)
Where $\alpha$ controls the strength of the balanced consistency term, $\lambda$ is a small positive coefficient of the entropy regularizer:
$\mathcal{H}(p(\mathbf{x}))=-\sum_{i}p_{i}(\mathbf{x})\log p_{i}(\mathbf{x})$ (6)
which is calculated based on the routing distribution $p(\mathbf{x})$ of each token. The first term enforces that routing intentions are consistent with actual usage, while the second term penalizes excessively high entropy in expert selection, thereby promoting more specialized expert assignments.
This combination ensures that the router not only maintains overall balance, but also forms a preference structure that avoids the problem that every expert is used equally regardless of the input semantics. We have analyzed RSL and traditional auxiliary loss from a theoretical perspective. Please refer to Appendix C for detailed derivation. For the balance loss in training, please refer to the Appendix  D.
Routing Optimization. After we prepare all LoRAs, we apply a soft training process on router. To prevent the expert knowledge in the first phase from being contaminated, we introduce a regularization term to penalize deviations from the previously learned expert parameters. Let the first phase parameters of expert $i$ be $\theta_{i}^{(0)}$ and the current parameters be $\theta_{i}$. We define the regularization term as:
$\mathcal{L}_{\text{preserve}}=\beta\cdot\sum_{i\in\mathcal{C}}\left\|\theta_{i}-\theta_{i}^{(0)}\right\|^{2}=\beta\cdot\sum_{i\in\mathcal{C}}\left\|\Delta\theta_{i}\right\|^{2},$ (7)
where $\mathcal{C}$ represents the set of constrained experts and $\beta$ controls the regularization strength. This regularization term constrains the sensitive experts to stay close to their original knowledge while still allowing other experts to adjust. For complex tasks, we support expert-level control, thus enabling flexible multi-expert learning.
To ensure that all experts can obtain meaningful gradients during the joint training process and promote stable optimization of the routing balance loss, we adopt soft expert fusion in training. Specifically, the router outputs the softmax routing scores $\mathbf{p}_{b,t}\in\mathbb{R}^{K}$ of all experts and fuses them, thereby achieving a fully differentiable mixture of LoRA experts. Although soft routing provides stable optimization and gradient propagation for all experts, its combination with the auxiliary loss leads to the problem of extremely balanced expert usage, that is, all experts are activated to the same extent regardless of the input semantics. To address this limitation, we introduce the Route-Specialization Balancing Loss (RSL) 3.3, which promotes the specialization of experts without sacrificing load balancing. The total loss during joint training becomes:
$\mathcal{L}_{\text{total}}=\underbrace{\mathcal{L}_{\text{task}}}_{\text{Supervised Objective}}+\underbrace{\alpha\cdot\mathcal{L}_{\text{RSL}}}_{\text{Routing Specialization}}+\underbrace{\beta\cdot\sum_{i\in\mathcal{C}}\left\|\theta_{i}-\theta_{i}^{(0)}\right\|^{2}}_{\text{Expert Preservation}},$ (8)
where $\mathcal{L}_{\text{task}}$ denotes the standard task loss (e.g., Cross-Entropy Loss). $\alpha$ is weighting factor for RSL. $\mathcal{L}_{\text{preserve}}$ is the expert regularization loss described above, scaled by $\beta$.
At inference time, we adopt a Top-3 routing strategy. By separating soft fusion at training time from sparse activation at inference time, we can ensure that the model obtains robust gradient propagation during routing optimization while performing efficient inference.
4 Experiment
To evaluate the effectiveness of our proposed LoRA-Mixer framework, we conduct experiments with LoRAs finetuned in five domains: Medical QA, Commonsense Reasoning, Natural Language Understanding, Mathematical Reasoning, and Coding Ability.
4.1 Experimental Setup
Datasets. To evaluate LoRA-Mixer, we selected seven publicly available benchmarks. For medical question answering, we used the Medical-QA dataset. For commonsense reasoning, we adopted ARC-E [30] and ARC-C [30], both of which focus on multiple-choice questions. For natural language understanding, we used SST2 [31] and CoLA [31], where SST2 is used for sentiment classification and CoLA is used for grammatical judgment. For mathematical reasoning, we chose the GSM8K [29] dataset, which contains thousands of elementary school math problems that require multi-step solutions. Finally, to evaluate encoding ability, we chose the HumanEva [19]l dataset.
Baselines. We chose three open source LLMs - LLaMA3-8B [26], Mistral-7B [27] and Falcon-Mamba-7B [28]. LLaMA3-8B and Mistral-7B both rely on the Transformer backbone network; Falcon-Mamba-7B is built entirely on the Mamba paradigm. In addition, we compared LoRA-Mixer with other state-of-the-art methods, including MoLE [14],MixLoRA [13],LoraHub [15],LoRA-LEGO [16].
Evaluation metrics We use three evaluation metrics to measure the performance. Specifically, for GSM8K [29], ARC-E [30], ARC-C [30], SST2 [31], and CoLA [31], we use ACC to measure performance. For the HumanEval dataset [19], we adopt the $Pass@1$ metric, which represents the ability of a single generated answer to correctly solve the task. Finally, considering the domain-specific freedom and rigor required by the Medical-QA dataset, we use DeepSeek-R1 [22] to evaluate completeness, correctness, and logical clarity, and report the final percentage scores.
4.2 Comparisons
Table1 shows the performance indicators of the three basemodels on various tasks. To ensure the reliability of the experimental results, we ran all experiments three times and took the average value.
Table 1: Performance of three base models, Falcon-Mamba-7B, Mistral-7B, and LLaMA3-8B, on seven benchmarks.
Base Model Medical CoLA SST2 GSM8K ARC-E ARC-C HumanEval
Falcon-Mamba-7B[28] 73.67 82.42 92.81 52.54 77.61 68.78 29.29
Mistral-7B[27] 66.32 71.21 85.24 40.83 80.00 61.50 27.95
LLaMA3-8B[26] 78.47 79.14 93.12 57.92 88.45 78.65 52.44
We compare our method with the state-of-the-art methods [14, 13, 15, 16] on seven datasets, and the experimental results are shown in Table 2. As can be seen from Table 2, our method achieves better performance than the baseline on most tasks. For the Falcon-Mamba , our method significantly outperforms other baselines on all tasks. For model details, please refer to Appendix B.
Table 2: Comparison of our LoRA-Mixer with LoRAHub [15], MoLE [14], and Mix-LoRA [13] across seven tasks (best scores in bold). Note that MixLoRA is excluded from Falcon-Mamba due to its Transformer-specific design.
Metho Medical CoLA SST2 GSM8K ARC-E ARC-C HumanEval
Falcon-Mamba (7B)
LoRAHub [15] 70.14 81.11 93.35 51.64 81.16 72.37 30.68
MOLE [14] 74.51 84.77 94.22 54.28 83.46 76.61 33.57
LoRA 77.26 85.62 95.07 56.27 85.68 76.51 33.54
LoRA-Mixer (ours) 78.01 85.91 95.76 57.87 86.87 77.19 35.37
Mistral (7B)
LoRAHub [15] 69.17 75.73 90.21 44.94 81.14 69.21 32.60
MOLE [14] 71.07 78.51 94.17 45.31 85.68 68.77 35.37
MixLoRA [13] 69.74 78.61 93.44 45.50 85.42 69.15 33.80
LoRA 70.33 79.19 93.58 46.67 86.66 70.53 35.31
LoRA-Mixer (ours) 71.25 82.17 95.16 46.48 87.87 71.22 36.76
LLaMA-3 (8B)
LoRAHub [15] 78.11 79.84 92.77 59.10 87.13 80.14 52.83
MOLE [14] 78.43 81.37 94.18 63.81 88.15 81.77 55.87
MixLoRA [13] 79.87 80.67 94.22 64.44 88.70 82.90 55.49
LoRA 81.09 81.50 95.30 65.14 89.59 82.15 55.61
LoRA-Mixer(ours) 81.55 82.22 95.41 65.53 89.88 83.24 57.32
Table 3: Evaluation of LoRA-Mixer on LoRAs sourced from Internet on five GLUE tasks, the base model is Flan-T5[36].
Method SST-2 CoLA MRPC RTE QQP
Flan-T5 [36] 94.01 74.21 79.90 80.08 82.32
LoRA 94.50 80.54 83.76 83.47 85.55
LoRA-Mixer 95.07 82.14 85.15 85.31 84.75
Table 4: Comparison of LoRA-Mixer and LoRA-LEGO. Results for LoRA-LEGO are from its paper [16].
Method CoLA SST-2 MRPC RTE
LoRA 61.63 75.74 68.00 52.22
LEGO [16] 55.48 73.22 66.00 71.85
LoRA-Mixer 64.60 80.31 72.24 61.47
Since LoRA-LEGO [16] is not open-sourced, to compare fairly, we use LLaMA2-7B [33] as the base model and conduct experiments on four tasks including CoLA, SST2, MRPC and RTE from LoRA-LEGO paper. Lora’s configuration employs a low rank of $r=6$ and a scaling factor of $\alpha=12$. The experimental results are shown in Table 4. From the results, it can be seen that our method outperforms LoRA-LEGO in three of the four tasks.
Considering that Mistral-7B and LLaMA3-8B have exactly the same model architecture, we directly migrate the parameters trained on Mistral-7B to LLaMA3-8B without any fine-tuning and adaptation, and conduct experiments on three datasets: ARC-E, ARC-C, and GSM8K. The results are shown in Table 5 . It is worth noting that we use the Zero-Shot CoT method to test the base model in the GSM8K task, and the results under different Few-Shot settings are also shown in Table 5.
Table 5: Evalution on LoRA-Mixer parameter transferability from Mistral-7B to LLaMA3-8B. Values show absolute performance (relative to baseline in parentheses).
Method GSM8K ARC-E ARC-C
0-shot 2-shot 5-shot 0-shot 0-shot
LLaMA3-8B 57.92 (1.00) 75.88 (1.00) 78.64 (1.00) 88.45 (1.00) 78.65 (1.00)
+ Mistral 59.13 (1.02) 76.26 (1.01) 81.43 (1.04) 85.89 (0.97) 79.14 (1.01)
Interestingly, we observed that we achieved better performance than the LLaMA3-8B baseline on two of the three tasks. This cross-model migration verifies that the LoRA expert and the learned routing function are not tightly coupled with a specific base model, making it possible to share experts between models with the same architecture.
4.3 Testing on LoRAs Sourced from Internet
To verify flexibility and plugin and play nature of LoRA-Mixer, we test our LoRA-Mixer on LoRAs sourced from Internet. Specifically, we download five distinct LoRAs from LoRAHub [15], which were trained on SST2, CoLA, MRPC, RTE, and QQP, respectively (please refer to Appendix E for more details). We use the Flan-T5 [36] model as a base model and mount the LoRAs without any modification, and collect $2K$ mixed data for routing training. The results are shown in Table 4. LoRA-Mixer achieved better performance on the four tasks, confirming that LoRA-Mixer’s potential for product-ready multitask applications.
4.4 Ablation Study
Impact of LoRA Rank. To evaluate the impact of rank in low-rank adaptation, we conducted experiments on $r=16$, $r=32$, $r=64$, and $r=128$, while keeping all other hyperparameters (e.g., dropout rate, learning rate) unchanged. The results of $r=64$ can be found in Table 2. We place the remaining results in Appendix A.
Expert Load Analysis. To analyze the overall load of experts, we uniformly sampled 1K data for seven benchmarks, including Medical, CoLA, SST2, GSM8K, ARC-E, ARC-C, and Humaneval. We report the average load of each expert on these $1K$ data, as shown in Figure 4. The activation rates of different experts are quite balanced, ranging from 15% - 18%, but in different tasks, the expert load reflects a kind of "perception" ability, and the expert load of specific tasks is higher than that of other experts, as shown in Figure  5. This shows that our routing mechanism effectively avoids expert collapse and achieves a balance of expert utilization between different tasks.
The impact of Top-$K$. To explore the effect of K values on Top-K routing, we used Falcon-Mamba as the basemodel to experiment on SST-2 and CoLA. The experimental results are shown in Figure  3.
Figure 3: Top-K Routing Impact.
Figure 4: Expert Assignment Overview.
As $K$ increases from 1 to 3, we observe that the accuracy of both tasks improves, indicating that using multiple experts allows the model to obtain complementary information. However, further increasing the value of K does not guarantee better results, but may degrade the performance. Therefore, the setting of $K$ is crucial for the MoE model. How to set or dynamically learn the most appropriate $K$ value is a direction worthy of further research in the future.
Table 6: The impact of LoRA-Mixer optimization on individual LoRAs (LoRA means adding independently trained LoRA, Expert means LoRA optimized by LoRA-Mixer).
Task LoRA w/o Expert w/ Expert Gap (Expert)
SST-2 95.30 94.70 95.41 +0.71
CoLA 81.50 80.15 82.22 +2.07
GSM8K 65.14 60.17 65.53 +5.36
HumanEval 55.61 53.39 57.32 +3.93
Enhanced Expert Performance Analysis. In order to verify the enhanced performances of individual expert after LoRA-Mixer optimization. We selected four tasks, GSM8K, SST2, CoLA and HumanEval, for experiments. The results are shown in Table 6. We can find that LoRAs optimized by LoRA-mixer exhibit improved performance, especially in the GSM8K task, where the performance improved by 5.36% after adding mathematical experts. This result confirms the enhanced ability of each individual LoRA expert after LoRA-Mixer optimization.
Cross-domain QA To evaluate the cross-domain generalization ability of LoRA-Mixer, we constructed two question-answering datasets: Medical-Mathematics and Mathematics-Coding. Each dataset contains 200 samples generated by DeepSeek-R1. These questions are more challenging. In the Medical-Mathematics dataset, the model needs to provide effective medical advice and corresponding calculations. In the Mathematics-Coding dataset, the model needs to generate correct Python code based on mathematical problems. The evaluation results of DeepSeek-R1 are shown in Table  7.
Table 7: Cross-domain performance of LoRA-Mixer on LLaMA3-8B [26].
Task Base LoRAHub [15] MOLE [14] MixLoRA [13] LoRA-Mixer
Math-Medical 69.88 70.53 72.11 72.74 73.41
Math-Coding 59.37 61.08 62.24 63.10 63.46
The impact of the RSL. We study the impact of our proposed RSL loss function on the LoRA-Mixer framework. RSL has two major advantages.
First, it enables routers to achieve global load balancing while maintaining strong input perception, which is a key factor to fully exploit the potential of sparse expert models. Second, compared with traditional load balancing loss functions, RSL significantly reduces the amount of training data required for effective router optimization.
To verify the first conclusion, we conducted experiments on three tasks: Medical, GSM8K, and HumanEval. As shown in Figure 5, after using RSL loss, LoRA-Mixer always assigns higher activation weights to experts related to the target task, which reflects the router’s strong domain perception and adaptive specialization capabilities. In contrast, using only auxiliary loss, the router will evenly distribute experts regardless of the semantics of the input content, which will result in the potential of the relevant experts not being fully developed, causing a performance bottleneck.
To support the second conclusion, we analyze the impact of training data size on routing performance. Specifically, we construct training sets of different sizes by sampling from a multi-task dataset pool and evaluate the performance on seven benchmark tasks. For clarity, we report the average performance over all tasks.
Table 8: Average performance across seven tasks under different routing training data sizes, with or without RSL. With RSL, LoRA-Mixer requires much less data while showing better performances.
Data w/ RSL w/o RSL Gap
1K 76.80 75.47 +1.33
2K 79.26 77.29 +1.97
3K 78.64 77.54 +1.10
4K 78.51 79.28 -0.77
Figure 5: Expert Load Distribution across Tasks.
As shown in Table 4.4, LoRA-Mixer achieves comparable performance using only 51.62% of the routing supervision data required by traditional auxiliary loss methods. We attribute this data efficiency to the dual regularization mechanism introduced by RSL. Specifically, a global consistency term $\bar{\mathbf{p}}^{\top}\bar{\mathbf{f}}$ aligns the expected routing probabilities with actual expert utilization, while a local token-level entropy penalty encourages diverse and selective expert activation. This synergistic design mitigates the overly uniform expert usage common in auxiliary loss, promoting both expert specialization and routing sparsity. As a result, the model maintains robust and adaptive expert assignment even under limited supervision. A detailed theoretical justification is provided in Appendix C.
5 Conclusion and Discussion
This paper introduces LoRA-Mixer, a flexible and architecture-agnostic MoE framework for combining LoRAs, adapting LLM for multitask. It improves the performance of Transformer and SSM models by replacing the projection layer with a dynamically routed LoRA experts. Through a two-stage training paradigm, LoRA-Mixer decouples expert learning from routing, enabling specialization and task awareness. To address the problem of overly uniform auxiliary losses, we propose RSL, which balances expert load while improving routing selectivity. The framework enables efficient router training with minimal data and supports cross-domain reuse of LoRA modules. Although LoRA-Mixer is effective, its fixed top-$K$ routing may limit adaptability to ambiguous inputs. Uniformly applying it across all layers can also introduce redundancy, as different layers capture different information. Future work will explore dynamic or differentiable routing and adaptive integration to apply LoRA-Mixer only where most beneficial.
Appendix A Experiment Result
Table 9: Comparison of LoRA-Mixer on Falcon-Mamba, Mistral, and LLaMA across seven tasks. LoRA denotes single-task fine-tuning with rank $r=16$. Best results per model and task are highlighted in bold.
Methods Medical CoLA SST2 GSM8K ARC-E ARC-C HumanEval
FalconMamba-LoRA 76.33 82.75 93.23 54.62 83.97 76.08 28.66
+LoRA-Mixer 77.03 83.80 93.41 55.15 84.17 76.51 29.48
Mistral-LoRA 67.87 75.55 89.14 45.96 84.37 69.51 34.76
+LoRA-Mixer 68.27 77.64 90.27 45.61 84.46 70.15 34.68
LLaMA-LoRA 79.35 77.65 94.15 61.79 88.64 79.47 51.78
+LoRA-Mixer 79.88 78.11 94.97 61.14 89.29 79.87 53.39
Table 10: Comparison of LoRA-Mixer on Falcon-Mamba, Mistral, and LLaMA across seven tasks. LoRA denotes single-task fine-tuning with rank $r=32$. Best results per model and task are highlighted in bold.
Methods Medical CoLA SST2 GSM8K ARC-E ARC-C HumanEval
Falcon-Mamba-LoRA 76.32 85.90 93.12 54.76 84.86 75.67 32.33
+LoRA-Mixer 76.67 86.00 95.35 55.41 85.55 76.81 34.15
Mistral-LoRA 68.57 76.89 93.87 46.29 84.87 68.83 32.93
+LoRA-Mixer 68.88 78.81 94.60 45.91 85.91 71.80 33.56
LLaMA-LoRA 79.15 81.11 95.30 61.38 88.76 79.31 52.34
+LoRA-Mixer 80.87 81.30 95.53 62.46 89.04 79.48 53.65
Table 11: Comparison of LoRA-Mixer on Falcon-Mamba, Mistral, and LLaMA across seven tasks. LoRA denotes single-task fine-tuning with rank $r=128$. Best results per model and task are highlighted in bold.
Methods Medical CoLA SST2 GSM8K ARC-E ARC-C HumanEval
Falcon-Mamba-LoRA 76.42 85.25 92.88 55.25 83.91 75.72 32.41
+LoRA-Mixer 76.81 85.97 94.30 56.11 85.50 77.75 33.10
Mistral-LoRA 69.63 79.85 90.14 46.05 84.62 68.59 34.87
+LoRA-Mixer 69.82 80.75 91.55 44.55 85.85 71.74 35.17
LLaMA-LoRA 79.38 81.48 95.27 62.04 89.21 80.28 55.40
+LoRA-Mixer 80.82 82.25 95.50 63.41 89.33 81.43 56.31
As shown in Table LABEL:r16,Table 10 and Table  11, within a certain range, as $r$ increases, the performance of the model can be improved to a certain extent. Our method is not only better than the basic model, but even better than the fine-tuned basic model on most tasks. This shows that our method can effectively combine existing knowledge through dynamic expert combination to form a more "intelligent" model.
Appendix B Falcon-Mamba Architecture Analysis
Mamba builds upon the state space model. It processes an input sequence $x(t)\in\mathbb{R}^{L}$ to produce an output $y(t)\in\mathbb{R}^{L}$ by employing a hidden state $h(t)\in\mathbb{R}^{N}$. This relationship is initially defined by a continuous system:
$h^{\prime}(t)$ $=\mathbf{A}h(t)+\mathbf{B}x(t),$ (9)
$y(t)$ $=\mathbf{C}h(t).$
Here, $\mathbf{A}\in\mathbb{R}^{N\times N}$ is the state transition matrix, and $\mathbf{B}\in\mathbb{R}^{N\times 1}$, $\mathbf{C}\in\mathbb{R}^{N\times 1}$ are projection matrices.
To process discrete sequences, Mamba discretizes the continuous parameters $\mathbf{A}$ and $\mathbf{B}$ using a time scale parameter $\Delta$ and the zero-order hold (ZOH) principle, resulting in discretized parameters $\mathbf{\overline{A}}$ and $\mathbf{\overline{B}}$:
$\mathbf{\overline{A}}$ $=\exp\left(\mathbf{\Delta A}\right),$ (10)
$\mathbf{\overline{B}}$ $=\left(\mathbf{\Delta A}\right)^{-1}\left(\exp\left(\mathbf{\Delta A}\right)-\mathbf{I}\right)\cdot\mathbf{\Delta B}.$
The discrete state-space equation with a step size of $\Delta$ becomes:
$h_{t}$ $=\mathbf{\overline{A}}h_{t-1}+\mathbf{\overline{B}}x_{t},$ (11)
$y_{t}$ $=\mathbf{C}h_{t}.$
By iteratively expanding the hidden state $h_{t-1}$, Mamba derives a global convolution kernel $\mathbf{\overline{K}}\in\mathbb{R}^{L}$. This kernel is then used to compute the output $y$ through a convolution operation with the input $x$:
$\mathbf{\overline{K}}$ $=\left(\mathbf{C\overline{B},C\overline{A}\overline{B},...,C\overline{A}^{\mathrm{L}-1}\overline{B}}\right),$ (12)
$y$ $=x\otimes\mathbf{\overline{K}}.$
Falcon Mamba 7B adopts a pure Mamba architecture, a departure from hybrid designs incorporating staggered attention. This deliberate choice aims to maintain the intrinsic linear scalability characteristic of Mamba models. To enhance adaptability, the model employs decoupled input embeddings and output weights.
At its core, Falcon-Mamba features 64 layers of the Falcon-Mamba Mixer. Each Mixer layer integrates an SSM (State Space Model) module alongside in-projection and out-projection layers, RMS Norm, and a convolutional layer.
Within the SSM module, the input is mapped to $\Delta$, $B$, and $C$ through a projection layer denoted as $x$-proj:
$x\xrightarrow{x\text{-proj}}(\Delta,B,C)$
where $x$ represents the input to the SSM module. Furthermore, another projection layer, $dt$-proj, discretizes $\Delta$:
$\Delta\xrightarrow{dt\text{-proj}}\Delta_{discretized}$
These discretized values—$\Delta_{discretized}$, $A$, $B$, $C$, and $D$—are then fed into the selective scanning module for processing. This architectural design of Falcon-Mamba fully enables the application of LoRA-Mixer specifically tailored for the projection layer. For a comprehensive understanding of the training process, please refer to.
Appendix C Theoretical Justification of RSL Loss
We provide a theoretical analysis of the proposed RSL loss and contrast it with the conventional auxiliary loss. Our goal is to demonstrate that RSL naturally promotes input-aware, expert-specialized routing with improved data efficiency.
C.1 Preliminaries
Let the router output a softmax distribution $p(x)=[p_{1}(x),\dots,p_{K}(x)]\in\Delta^{K-1}$ over $K$ experts for a token $x$, where
$p_{i}(x)=\frac{\exp(G_{i}(x))}{\sum_{j=1}^{K}\exp(G_{j}(x))},\quad\sum_{i=1}^{K}p_{i}(x)=1.$ (13)
We define the expected routing probability and top-1 selection frequency as:
$\bar{p}_{i}$ $=\mathbb{E}_{x\sim\mathcal{D}}[p_{i}(x)],$ (14)
$\bar{f}_{i}$ $=\mathbb{E}_{x\sim\mathcal{D}}\left[\mathbb{I}(i=\arg\max_{j}p_{j}(x))\right],$ (15)
where $\bar{p}_{i}$ represents the average routing intention, and $\bar{f}_{i}$ represents the empirical usage under hard top-1 routing.
C.2 Auxiliary Loss and Its Implicit Bias
The standard auxiliary loss encourages load balancing by aligning the average routing with actual expert usage:
$\mathcal{L}_{\text{aux}}=\alpha\sum_{i=1}^{K}\bar{p}_{i}\cdot\bar{f}_{i}.$ (16)
Proposition 1 (Equilibrium of Auxiliary Loss).
Under the constraint $\sum_{i=1}^{K}\bar{p}_{i}=1$, the minimum of $\mathcal{L}_{\text{aux}}$ is attained when $\bar{p}_{i}=\bar{f}_{i}$ for all $i$.
Proof. Using the Lagrangian method:
$\mathcal{L}=\sum_{i=1}^{K}\bar{p}_{i}\bar{f}_{i}-\lambda\left(\sum_{i=1}^{K}\bar{p}_{i}-1\right)$
Taking partial derivatives:
$\frac{\partial\mathcal{L}}{\partial\bar{p}_{i}}=\bar{f}_{i}-\lambda=0\Rightarrow\bar{f}_{i}=\lambda,\quad\forall i.$
Thus, all $\bar{f}_{i}$ are equal, implying uniform distribution: $\bar{f}_{i}=\frac{1}{K}$, hence $\bar{p}_{i}=\frac{1}{K}$.
This shows that the auxiliary loss alone biases the router toward uniform expert activation, regardless of input characteristics.
C.3 RSL Loss: Entropy-Regularized Routing
To promote more input-sensitive routing, we introduce an entropy-regularized objective:
$\mathcal{L}_{\text{RSL}}=\mathcal{L}_{\text{aux}}-\lambda\cdot\mathbb{E}_{x\sim\mathcal{D}}[\mathcal{H}(p(x))],$ (17)
where the token-level entropy is defined as:
$\mathcal{H}(p(x))=-\sum_{i=1}^{K}p_{i}(x)\log p_{i}(x).$ (18)
This term encourages the router to assign higher weights to fewer experts, effectively breaking uniformity and encouraging selective specialization.
C.4 Gradient Analysis and Token-Awareness
We derive the entropy gradient w.r.t. routing score $p_{i}(x)$:
$\frac{\partial\mathcal{H}(p(x))}{\partial p_{i}(x)}$ $=-\log p_{i}(x)-1,\quad\text{subject to}\quad\sum_{i=1}^{K}p_{i}(x)=1,$ (19)
$=-\log p_{i}(x)-1+\mu,$ (20)
where $\mu$ is the Lagrange multiplier due to the simplex constraint.
Thus, the total gradient of the RSL loss becomes:
$\nabla_{p_{i}(x)}\mathcal{L}_{\text{RSL}}=\alpha\cdot\frac{\partial\bar{p}_{i}}{\partial p_{i}(x)}\cdot\bar{f}_{i}+\lambda(\log p_{i}(x)+1-\mu).$ (21)
This shows that RSL introduces token-level gradient signals via $\log p_{i}(x)$, unlike the auxiliary loss, which propagates only global gradients.
C.5 Token-Awareness via Routing Variance
To quantify input-aware routing, we define:
$\text{Var}_{x\sim\mathcal{D}}(p(x)):=\mathbb{E}_{x}\left[\|p(x)-\bar{p}\|^{2}\right].$ (22)
We say the routing is token-aware if $\text{Var}(p(x))>\epsilon$ for some $\epsilon>0$. The auxiliary loss tends to reduce this variance (driving uniform routing), while RSL encourages high variance and peaked distributions aligned with input semantics.
C.6 Conclusion
The RSL loss incorporates an entropy-based regularizer that mitigates the uniformity bias of auxiliary loss. By injecting token-level gradient signals and promoting routing variance, RSL enables input-aware, specialized, and discriminative expert assignments. This property is especially beneficial in data-scarce regimes, where each token’s contribution to routing must be maximally leveraged. Note that RSL is fully compatible with auxiliary loss; it can be viewed as a strict generalization that stabilizes training while encouraging specialization.
Figure 6: Balance loss curve using RSL loss during training.
Appendix D Balance loss visualization
Figure 6 shows the changing trend of the Balance Loss when the RSL loss function is used during training. As shown in the figure, the Balance Loss drops rapidly in the early stage of training, indicating that our model can quickly learn an effective expert routing strategy, thanks to the synergy of the global consistency term and the local entropy penalty term in the RSL loss function. In the middle stage of training, the Balance Loss continues to drop steadily with a small fluctuation, which reflects the stability that the RSL loss function brings to the training process. In the late stage of training, the Balance Loss remains stable at a low level, further demonstrating the balance and optimization effect of the model in the use of experts. Overall, this Balance Loss curve not only reflects the model’s ability to converge quickly, but also demonstrates the robustness of the training process, verifying the significant advantages of the RSL loss function in improving model performance and training efficiency.
Appendix E Details of LoRA modules of Flan-T5.
Parameter Value
base_model_name_or_path google/flan-t5-large
bias none
fan_in_fan_out false
inference_mode true
init_lora_weights true
layers_pattern null
layers_to_transform null
lora_alpha 32
lora_dropout 0.1
modules_to_save null
peft_type LORA
r 16
revision null
target_modules [q, v]
task_type SEQ_2_SEQ_LM
Table 12: LoRA configuration details used in our experiments.
Appendix F Experimental details
Our experiments are conducted on a Linux workstation equipped with a single NVIDIA A800 80GB GPU and a 32-core Intel Xeon CPU. We use the AdamW optimizer with a learning rate of $1\times 10^{-5}$. For Transformer-based models, LoRA-Mixer is applied exclusively to the attention modules. For SSM-based models, LoRA-Mixer is integrated into the in, out, dt, and x projection layers.
NeurIPS Paper Checklist
  1. 1.
    Claims
  2. Question: Do the main claims made in the abstract and introduction accurately reflect the paper’s contributions and scope?
  3. Answer: [Yes]
  4. Justification: The abstract and introduction of the paper accurately reflect the contributions of the paper, such as proposing LoRA-Mixer, a highly flexible MoE framework, and verifying its effectiveness through a large number of experiments.
  5. Guidelines:
    • The answer NA means that the abstract and introduction do not include the claims made in the paper.
    • The abstract and/or introduction should clearly state the claims made, including the contributions made in the paper and important assumptions and limitations. A No or NA answer to this question will not be perceived well by the reviewers.
    • The claims made should match theoretical and experimental results, and reflect how much the results can be expected to generalize to other settings.
    • It is fine to include aspirational goals as motivation as long as it is clear that these goals are not attained by the paper.
  6. 2.
    Limitations
  7. Question: Does the paper discuss the limitations of the work performed by the authors?
  8. Answer: [Yes]
  9. Justification: Due to the length limitations of the article, we did not include a separate section in the main text to discuss the limitations of the article. We briefly discuss the limitations in the Appendix.
  10. Guidelines:
    • The answer NA means that the paper has no limitation while the answer No means that the paper has limitations, but those are not discussed in the paper.
    • The authors are encouraged to create a separate "Limitations" section in their paper.
    • The paper should point out any strong assumptions and how robust the results are to violations of these assumptions (e.g., independence assumptions, noiseless settings, model well-specification, asymptotic approximations only holding locally). The authors should reflect on how these assumptions might be violated in practice and what the implications would be.
    • The authors should reflect on the scope of the claims made, e.g., if the approach was only tested on a few datasets or with a few runs. In general, empirical results often depend on implicit assumptions, which should be articulated.
    • The authors should reflect on the factors that influence the performance of the approach. For example, a facial recognition algorithm may perform poorly when image resolution is low or images are taken in low lighting. Or a speech-to-text system might not be used reliably to provide closed captions for online lectures because it fails to handle technical jargon.
    • The authors should discuss the computational efficiency of the proposed algorithms and how they scale with dataset size.
    • If applicable, the authors should discuss possible limitations of their approach to address problems of privacy and fairness.
    • While the authors might fear that complete honesty about limitations might be used by reviewers as grounds for rejection, a worse outcome might be that reviewers discover limitations that aren’t acknowledged in the paper. The authors should use their best judgment and recognize that individual actions in favor of transparency play an important role in developing norms that preserve the integrity of the community. Reviewers will be specifically instructed to not penalize honesty concerning limitations.
  11. 3.
    Theory assumptions and proofs
  12. Question: For each theoretical result, does the paper provide the full set of assumptions and a complete (and correct) proof?
  13. Answer: [Yes]
  14. Justification: In the appendix, we provide a thorough theoretical analysis on why the RSL function is more effective than traditional auxiliary loss.
  15. Guidelines:
    • The answer NA means that the paper does not include theoretical results.
    • All the theorems, formulas, and proofs in the paper should be numbered and cross-referenced.
    • All assumptions should be clearly stated or referenced in the statement of any theorems.
    • The proofs can either appear in the main paper or the supplemental material, but if they appear in the supplemental material, the authors are encouraged to provide a short proof sketch to provide intuition.
    • Inversely, any informal proof provided in the core of the paper should be complemented by formal proofs provided in appendix or supplemental material.
    • Theorems and Lemmas that the proof relies upon should be properly referenced.
  16. 4.
    Experimental result reproducibility
  17. Question: Does the paper fully disclose all the information needed to reproduce the main experimental results of the paper to the extent that it affects the main claims and/or conclusions of the paper (regardless of whether the code and data are provided or not)?
  18. Answer: [Yes]
  19. Justification: The experimental section of our article fully describes the experimental configuration we used, such as the configuration of LoRA, the details of the dataset, the evaluation criteria used, etc. To ensure the reproducibility of the article, we will open source the code.
  20. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • If the paper includes experiments, a No answer to this question will not be perceived well by the reviewers: Making the paper reproducible is important, regardless of whether the code and data are provided or not.
    • If the contribution is a dataset and/or model, the authors should describe the steps taken to make their results reproducible or verifiable.
    • Depending on the contribution, reproducibility can be accomplished in various ways. For example, if the contribution is a novel architecture, describing the architecture fully might suffice, or if the contribution is a specific model and empirical evaluation, it may be necessary to either make it possible for others to replicate the model with the same dataset, or provide access to the model. In general. releasing code and data is often one good way to accomplish this, but reproducibility can also be provided via detailed instructions for how to replicate the results, access to a hosted model (e.g., in the case of a large language model), releasing of a model checkpoint, or other means that are appropriate to the research performed.
    • While NeurIPS does not require releasing code, the conference does require all submissions to provide some reasonable avenue for reproducibility, which may depend on the nature of the contribution. For example
      1. (a)
        If the contribution is primarily a new algorithm, the paper should make it clear how to reproduce that algorithm.
      2. (b)
        If the contribution is primarily a new model architecture, the paper should describe the architecture clearly and fully.
      3. (c)
        If the contribution is a new model (e.g., a large language model), then there should either be a way to access this model for reproducing the results or a way to reproduce the model (e.g., with an open-source dataset or instructions for how to construct the dataset).
      4. (d)
        We recognize that reproducibility may be tricky in some cases, in which case authors are welcome to describe the particular way they provide for reproducibility. In the case of closed-source models, it may be that access to the model is limited in some way (e.g., to registered users), but it should be possible for other researchers to have some path to reproducing or verifying the results.
  21. 5.
    Open access to data and code
  22. Question: Does the paper provide open access to the data and code, with sufficient instructions to faithfully reproduce the main experimental results, as described in supplemental material?
  23. Answer: [Yes]
  24. Justification: We will open source the code and submit it in the supplementary materials. As for the datasets, we have clearly stated them in the article. These datasets are open source and can fully guarantee the authenticity of the content of the article.
  25. Guidelines:
    • The answer NA means that paper does not include experiments requiring code.
    • Please see the NeurIPS code and data submission guidelines (https://nips.cc/public/guides/CodeSubmissionPolicy) for more details.
    • While we encourage the release of code and data, we understand that this might not be possible, so “No” is an acceptable answer. Papers cannot be rejected simply for not including code, unless this is central to the contribution (e.g., for a new open-source benchmark).
    • The instructions should contain the exact command and environment needed to run to reproduce the results. See the NeurIPS code and data submission guidelines (https://nips.cc/public/guides/CodeSubmissionPolicy) for more details.
    • The authors should provide instructions on data access and preparation, including how to access the raw data, preprocessed data, intermediate data, and generated data, etc.
    • The authors should provide scripts to reproduce all experimental results for the new proposed method and baselines. If only a subset of experiments are reproducible, they should state which ones are omitted from the script and why.
    • At submission time, to preserve anonymity, the authors should release anonymized versions (if applicable).
    • Providing as much information as possible in supplemental material (appended to the paper) is recommended, but including URLs to data and code is permitted.
  26. 6.
    Experimental setting/details
  27. Question: Does the paper specify all the training and test details (e.g., data splits, hyperparameters, how they were chosen, type of optimizer, etc.) necessary to understand the results?
  28. Answer: [No]
  29. Justification: Due to space limitations, we did not explore these details in detail in this paper.
  30. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • The experimental setting should be presented in the core of the paper to a level of detail that is necessary to appreciate the results and make sense of them.
    • The full details can be provided either with the code, in appendix, or as supplemental material.
  31. 7.
    Experiment statistical significance
  32. Question: Does the paper report error bars suitably and correctly defined or other appropriate information about the statistical significance of the experiments?
  33. Answer: [Yes]
  34. Justification: In order to reduce the error, the experimental results of our article are the average results obtained by repeating the experiment more than three times.
  35. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • The authors should answer "Yes" if the results are accompanied by error bars, confidence intervals, or statistical significance tests, at least for the experiments that support the main claims of the paper.
    • The factors of variability that the error bars are capturing should be clearly stated (for example, train/test split, initialization, random drawing of some parameter, or overall run with given experimental conditions).
    • The method for calculating the error bars should be explained (closed form formula, call to a library function, bootstrap, etc.)
    • The assumptions made should be given (e.g., Normally distributed errors).
    • It should be clear whether the error bar is the standard deviation or the standard error of the mean.
    • It is OK to report 1-sigma error bars, but one should state it. The authors should preferably report a 2-sigma error bar than state that they have a 96% CI, if the hypothesis of Normality of errors is not verified.
    • For asymmetric distributions, the authors should be careful not to show in tables or figures symmetric error bars that would yield results that are out of range (e.g. negative error rates).
    • If error bars are reported in tables or plots, The authors should explain in the text how they were calculated and reference the corresponding figures or tables in the text.
  36. 8.
    Experiments compute resources
  37. Question: For each experiment, does the paper provide sufficient information on the computer resources (type of compute workers, memory, time of execution) needed to reproduce the experiments?
  38. Answer: [No]
  39. Justification: Due to the length of the article, we do not discuss the computing resources in detail. We can briefly explain here that we use an A800 80G GPU for training and inference.
  40. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • The paper should indicate the type of compute workers CPU or GPU, internal cluster, or cloud provider, including relevant memory and storage.
    • The paper should provide the amount of compute required for each of the individual experimental runs as well as estimate the total compute.
    • The paper should disclose whether the full research project required more compute than the experiments reported in the paper (e.g., preliminary or failed experiments that didn’t make it into the paper).
  41. 9.
    Code of ethics
  42. Question: Does the research conducted in the paper conform, in every respect, with the NeurIPS Code of Ethics https://neurips.cc/public/EthicsGuidelines?
  43. Answer: [Yes]
  44. Justification: Our paper complies with the ethical standards of the conference in all aspects and will not cause any adverse impact on society.
  45. Guidelines:
    • The answer NA means that the authors have not reviewed the NeurIPS Code of Ethics.
    • If the authors answer No, they should explain the special circumstances that require a deviation from the Code of Ethics.
    • The authors should make sure to preserve anonymity (e.g., if there is a special consideration due to laws or regulations in their jurisdiction).
  46. 10.
    Broader impacts
  47. Question: Does the paper discuss both potential positive societal impacts and negative societal impacts of the work performed?
  48. Answer: [No]
  49. Justification: Due to the length limitation of the article, we did not discuss this aspect, but our research will not cause any adverse impact on society.
  50. Guidelines:
    • The answer NA means that there is no societal impact of the work performed.
    • If the authors answer NA or No, they should explain why their work has no societal impact or why the paper does not address societal impact.
    • Examples of negative societal impacts include potential malicious or unintended uses (e.g., disinformation, generating fake profiles, surveillance), fairness considerations (e.g., deployment of technologies that could make decisions that unfairly impact specific groups), privacy considerations, and security considerations.
    • The conference expects that many papers will be foundational research and not tied to particular applications, let alone deployments. However, if there is a direct path to any negative applications, the authors should point it out. For example, it is legitimate to point out that an improvement in the quality of generative models could be used to generate deepfakes for disinformation. On the other hand, it is not needed to point out that a generic algorithm for optimizing neural networks could enable people to train models that generate Deepfakes faster.
    • The authors should consider possible harms that could arise when the technology is being used as intended and functioning correctly, harms that could arise when the technology is being used as intended but gives incorrect results, and harms following from (intentional or unintentional) misuse of the technology.
    • If there are negative societal impacts, the authors could also discuss possible mitigation strategies (e.g., gated release of models, providing defenses in addition to attacks, mechanisms for monitoring misuse, mechanisms to monitor how a system learns from feedback over time, improving the efficiency and accessibility of ML).
  51. 11.
    Safeguards
  52. Question: Does the paper describe safeguards that have been put in place for responsible release of data or models that have a high risk for misuse (e.g., pretrained language models, image generators, or scraped datasets)?
  53. Answer: [No]
  54. Justification: The datasets and models we use are open source and safe. There is no risk.
  55. Guidelines:
    • The answer NA means that the paper poses no such risks.
    • Released models that have a high risk for misuse or dual-use should be released with necessary safeguards to allow for controlled use of the model, for example by requiring that users adhere to usage guidelines or restrictions to access the model or implementing safety filters.
    • Datasets that have been scraped from the Internet could pose safety risks. The authors should describe how they avoided releasing unsafe images.
    • We recognize that providing effective safeguards is challenging, and many papers do not require this, but we encourage authors to take this into account and make a best faith effort.
  56. 12.
    Licenses for existing assets
  57. Question: Are the creators or original owners of assets (e.g., code, data, models), used in the paper, properly credited and are the license and terms of use explicitly mentioned and properly respected?
  58. Answer: [Yes]
  59. Justification: The models and datasets we use are open source and have obtained asset licenses from their authors.
  60. Guidelines:
    • The answer NA means that the paper does not use existing assets.
    • The authors should cite the original paper that produced the code package or dataset.
    • The authors should state which version of the asset is used and, if possible, include a URL.
    • The name of the license (e.g., CC-BY 4.0) should be included for each asset.
    • For scraped data from a particular source (e.g., website), the copyright and terms of service of that source should be provided.
    • If assets are released, the license, copyright information, and terms of use in the package should be provided. For popular datasets, paperswithcode.com/datasets has curated licenses for some datasets. Their licensing guide can help determine the license of a dataset.
    • For existing datasets that are re-packaged, both the original license and the license of the derived asset (if it has changed) should be provided.
    • If this information is not available online, the authors are encouraged to reach out to the asset’s creators.
  61. 13.
    New assets
  62. Question: Are new assets introduced in the paper well documented and is the documentation provided alongside the assets?
  63. Answer: [No]
  64. Justification: We have not provided any explanation for this aspect and our model will also be open source.
  65. Guidelines:
    • The answer NA means that the paper does not release new assets.
    • Researchers should communicate the details of the dataset/code/model as part of their submissions via structured templates. This includes details about training, license, limitations, etc.
    • The paper should discuss whether and how consent was obtained from people whose asset is used.
    • At submission time, remember to anonymize your assets (if applicable). You can either create an anonymized URL or include an anonymized zip file.
  66. 14.
    Crowdsourcing and research with human subjects
  67. Question: For crowdsourcing experiments and research with human subjects, does the paper include the full text of instructions given to participants and screenshots, if applicable, as well as details about compensation (if any)?
  68. Answer: [No]
  69. Justification: Our experiments do not contain any experimental data related to humans and therefore need not be discussed.
  70. Guidelines:
    • The answer NA means that the paper does not involve crowdsourcing nor research with human subjects.
    • Including this information in the supplemental material is fine, but if the main contribution of the paper involves human subjects, then as much detail as possible should be included in the main paper.
    • According to the NeurIPS Code of Ethics, workers involved in data collection, curation, or other labor should be paid at least the minimum wage in the country of the data collector.
  71. 15.
    Institutional review board (IRB) approvals or equivalent for research with human subjects
  72. Question: Does the paper describe potential risks incurred by study participants, whether such risks were disclosed to the subjects, and whether Institutional Review Board (IRB) approvals (or an equivalent approval/review based on the requirements of your country or institution) were obtained?
  73. Answer: [No]
  74. Justification: Our experiments did not involve any experimental subjects.
  75. Guidelines:
    • The answer NA means that the paper does not involve crowdsourcing nor research with human subjects.
    • Depending on the country in which research is conducted, IRB approval (or equivalent) may be required for any human subjects research. If you obtained IRB approval, you should clearly state this in the paper.
    • We recognize that the procedures for this may vary significantly between institutions and locations, and we expect authors to adhere to the NeurIPS Code of Ethics and the guidelines for their institution.
    • For initial submissions, do not include any information that would break anonymity (if applicable), such as the institution conducting the review.
  76. 16.
    Declaration of LLM usage
  77. Question: Does the paper describe the usage of LLMs if it is an important, original, or non-standard component of the core methods in this research? Note that if the LLM is used only for writing, editing, or formatting purposes and does not impact the core methodology, scientific rigorousness, or originality of the research, declaration is not required.
  78. Answer: [Yes]
  79. Justification: The LLMs we use are all open source and We use the large language model for grammar polishing.
  80. Guidelines:
    • The answer NA means that the core method development in this research does not involve LLMs as any important, original, or non-standard components.
    • Please refer to our LLM policy (https://neurips.cc/Conferences/2025/LLM) for what should or should not be described.

### v2

@@
 Abstract
-Recent efforts to combine low-rank adaptation (LoRA) with mixture-of-experts (MoE) for adapting large language models (LLMs) to multiple tasks, yet exhibit prevailing limitations: they either swap entire attention/feed-forward layers for switch experts or bolt on parallel expert branches, diluting parameter efficiency and task fidelity.
-We propose the LoRA-Mixer, a modular and lightweight MoE framework that integrates LoRA experts.
-Our core innovation lies in replacing the projection matrices of the attention module’s input/output linear layers with dynamically routed, task-specific LoRA experts.
-This design ensures seamless compatibility with diverse foundation models—including transformers and state space models (SSMs)—by leveraging their inherent linear projection structures.
-The framework supports two operational paradigms: (1) joint optimization of LoRA experts and routing mechanisms via a novel hard-soft routing strategy, or (2) direct deployment of pre-trained, frozen LoRA modules sourced from external repositories.
-To enable robust router training with limited data while ensuring stable routing decisions and maximizing expert reuse, we introduce an adaptive Specialization Balance Loss (SBL) that jointly optimizes expert balance and task-specific alignment.
-Extensive experiments on seven benchmark datasets, including MedQA, CoLA, SST-2, GSM8K, ARC-E, ARC-C, and HumanEval, demonstrate the effectiveness of LoRA-Mixer.
-On datasets such as GSM8K, HumanEval, and MedQA, LoRA-Mixer achieves significant improvements of 7.61%, 4.88%, and 3.08% over the base models, respectively.
-Compared with the state-of-the-art methods, LoRA-Mixer achieves additional improvements of 1.09%, 1.45%, and 1.68%, respectively, using only 48% of the parameters, demonstrating its efficiency and strong performance.
-1 Introduction
-Large Language Models (LLMs) have achieved unprecedented proficiency in general-purpose reasoning and generation, yet their adaptation to specialized downstream domains remains computationally prohibitive, requiring significant resources for full-scale fine-tuning [1, 2].
-To mitigate these resource demands, parameter-efficient fine-tuning (PEFT) methods [37, 38, 39, 40, 41, 42] have emerged as a scalable paradigm for task-specific adaptation.
-Among these, Low-Rank Adaptation (LoRA) [4, 5] has demonstrated particular efficacy, operating through low-rank decomposition of updates to the pretrained weights—enabling efficient tuning with minimal parameter overhead.
-Recent work has explored modularly composing independently trained LoRA modules as a promising strategy for multitask adaptation; however, naive composition can result in interference between task-specific subspaces, limiting their synergistic potential[15, 14].
-This limitation has motivated exploration of mixture-of-experts (MoE) architecture [3, 6], which treat each task-specific LoRA as an expert and sparsely activate and fuse the experts.
-Recent studies demonstrate promising directions in hybrid Lo-
-Figure 1: MoE and LoRA-MoE integration methods.
-(1) and (2): replace the attention or feedforward layers with switch experts; (3) and (4): introduce LoRA experts branches in parallel with the attention or feedforward layers, and fuse the output into the main branch.
+Recent attempts to combine low-rank adaptation (LoRA) with mixture-of-experts (MoE) for multi-task adaptation of Large Language Models (LLMs) often replace whole attention/FFN layers with switch experts or append parallel expert branches, undermining parameter efficiency and limiting task specialization.
+We introduce LoRA-Mixer, a modular MoE framework that routes task-specific LoRA experts into the core projection matrices of the attention module (input/output linear layers), rather than primarily targeting FFN blocks.
+The design delivers fine-grained token-level specialization by fully exploiting the attention mechanism, while remaining drop-in compatible with Transformers and state-space models (SSMs) as the linear projection layers are ubiquitous.
+To train robust routers from limited data while promoting stable, selective decisions and high expert reuse, LoRA-Mixer employs an adaptive Routing Specialization Loss (RSL) that jointly enforces global load balance and input-aware specialization via an entropy-shaping objective.
+The framework supports two regimes: (i) joint optimization of adapters and router with a differentiable hard–soft top-$k$ routing scheme, and (ii) plug-and-play routing over frozen, pre-trained LoRA modules sourced from public repositories.
+Across 15 benchmarks—including MedQA, GSM8K, HumanEval, and GLUE—RSL-optimized LoRA-Mixer outperforms state-of-the-art routing and LoRA-MoE baselines while using 48% of their trainable parameters, with gains of +3.79%, +2.90%, and +3.95% on GSM8K, CoLA, and ARC-C, respectively.
+Cross-model transfer and adapter reuse experiments further demonstrate the approach’s versatility and data efficiency.
+Our code can be obtained at https://github.com/hustcselwb/LoRA-Mixer.
+1 Introduction.
+Large Language Models (LLMs) have achieved unprecedented proficiency in general-purpose reasoning and generation, yet their adaptation to specialized downstream domains remains computationally prohibitive for full-scale fine-tuning Brown et al.
+(2020); Touvron et al.
+(2023).
+To mitigate these resource demands, parameter-efficient fine-tuning (PEFT) methods Li & Liang (2021); Zaken et al.
+(2021); Liu et al.
+(2022); Houlsby et al.
+(2019); Liu et al.
+(2021) have emerged as a scalable paradigm.
+Among them, Low-Rank Adaptation (LoRA) Hu et al.
+(2022); Tian et al.
+(2024) has demonstrated particular efficacy, operating through low-rank decomposition of updates to the pretrained weights—enabling efficient tuning with minimal parameter overhead.
+Recent work has explored modularly composing independently trained LoRA modules as a promising strategy for multitask adaptation; however, naive composition can result in interference between task-specific subspaces, limiting their synergistic potential Huang et al.
+(2023); Wu et al.
+(2024).
+This limitation has motivated exploration of mixture-of-experts (MoE) architecture Shazeer et al.
+(2017); Fedus et al.
+(2022), which treat each task-specific LoRA as an expert and sparsely activate and fuse the experts.
+Recent studies demonstrate promising directions in hybrid LoRA-MoE frameworks Wu et al.
+(2024); Li et al.
+(2024a); Huang et al.
+(2023); Zhao et al.
+(2024); Liao et al.
+(2025); Gao et al.
+(2024); Chen et al.
+(2024); Dou et al.
+(2023), aiming to enhance model performance on complex tasks across multi-domain datasets while preserving the parameter efficiency of fine-tuning.
+Figure 1: LoRA and MoE integration methods.
+(1) and (2): replace the attention or FNNs with switch experts; (3) and (4): introduce LoRA experts branches in parallel with the attention or feedforward layers, and fuse the output into the main branch.
 Our LoRA-Mixer (right) applies mixture of LoRA experts to the projection layers, which can effectively leverages the attention mechanism.
-RA-MoE frameworks [14, 13, 15, 16, 35, 34, 23, 24], aiming to enhance model performance on complex tasks across multi-domain datasets while preserving the parameter efficiency of fine-tuning.
-The central challenge in composing multiple pre-trained LoRAs lies in effectively synergizing them—achieving enhanced performance across constituent tasks while minimizing training overhead and preserving their distinct, task-specific characteristics.
-Conventional methods integrating LoRA into mixture-of-experts (MoE) architectures typically follow one of two paradigms: (1) directly substituting attention or feedforward layers with LoRA-based switchable experts [13, 23], emulating the classical MoE structure [6, 3]; or (2) introducing parallel branches of LoRA experts whose outputs are subsequently fused into the primary model pipeline [14, 15], as illustrated in Fig.1.
-While these strategies have demonstrated promising empirical results [14, 35, 52, 53, 54, 55], they still encounter fundamental limitations.
-Specifically, the vanilla MoE-inspired paradigm necessitates joint training across all expert modules, significantly increasing training data demands and restricting the modular reuse and transferability of pre-trained LoRAs.
-Conversely, the parallel LoRA-expert approach circumvents the inherent attention or state transition mechanisms, resulting in simplistic output fusion and suboptimal overall integration.
-Additionally, auxiliary losses commonly employed for routing optimization inherently promote uniform load balancing among experts, diminishing the capacity for nuanced task-awareness [13].
-These limitations motivate the development of a more flexible, plug-and-play framework that is model-architecture agnostic, compatible with both Transformer and state-space model (SSM) architectures [7], capable of training an efficient and discriminative routing mechanism with minimal computational and data demands, and maximizing the reuse and transferability of independently pre-trained LoRA modules.
-In this paper, we introduce LoRA-Mixer, a novel framework designed to efficiently synergize multiple pre-trained LoRA modules by treating them as dynamic, pluggable memory cells.
-LoRA-Mixer equips the linear projection layers of the original model with mixed LoRA experts, enabling these experts to directly leverage the effectiveness of the core attention or state-transition mechanisms.
+The central challenge in composing pre-trained LoRAs is to realize synergy—improving performance across constituent tasks—without inflating training cost or erasing task-specific inductive biases.
+Existing LoRA–MoE integrations largely follow two patterns: (i) replacing attention or FFN blocks with LoRA-based switch experts in a standard MoE setup Li et al.
+(2024a); Chen et al.
+(2024); Fedus et al.
+(2022); Shazeer et al.
+(2017); or (ii) attaching parallel LoRA branches whose outputs are fused back into the main path Wu et al.
+(2024); Huang et al.
+(2023) (Fig.
+1).
+Despite encouraging results Wu et al.
+(2024); Liao et al.
+(2025); Krajewski et al.
+(2024); Zhang et al.
+(2019); Li et al.
+(2022); Gross et al.
+(2017), switch-style designs typically require joint training of all experts, increasing data demands and hindering modular reuse and transfer of off-the-shelf LoRAs.
+Parallel-branch schemes sidestep native attention or state-transition pathways, yielding shallow output fusion and weak integration.
+Moreover, common auxiliary routing losses emphasize uniform load balancing, suppressing input- and task-aware specialization Li et al.
+(2024a; 2026); Zhang et al.
+(2026c).
+These issues motivate a plug-and-play, architecture-agnostic alternative—compatible with both Transformers and state-space models (SSMs) Gu & Dao (2023); Zhang et al.
+(2026a; b)—that learns discriminative routing with minimal compute and data while maximizing reuse of independently trained LoRA modules.
+In response, we introduce LoRA-Mixer, a novel framework designed to efficiently synergize multiple pre-trained LoRA modules by treating them as dynamic, pluggable memory cells.
+LoRA-Mixer equips the linear projection layers of the original model with mixed LoRA experts, enabling these experts to directly leverage the core attention or state-transition mechanisms.
+To further enhance efficiency and maintain routing effectiveness, LoRA-Mixer adopts a novel Router Specialization Balancing Loss (RSL) to align routing decisions with token-level expert usage, maintaining moderate entropy to encourage exploratory behavior.
+During inference, we employ sparse top-$K$ fusion, effectively balancing computational cost and scalability without compromising expert selectivity.
 LoRA-Mixer supports LoRA modules sourced from external repositories, independently trained, or jointly trained through hard routing strategies, allowing seamless plug-and-play usage across various tasks and domains.
@@
 Consequently, LoRA-Mixer is particularly suitable for constructing large-scale, modular language models characterized by task-specific memory, computational efficiency, and strong transferability.
-To further enhance efficiency and maintain routing effectiveness, we propose a novel Router Specialization Balancing Loss (RSBL).
-RSL aligns routing decisions with token-level expert usage, maintaining moderate entropy to encourage exploratory behavior.
-During inference, we employ sparse top-$K$ fusion, effectively balancing computational cost and scalability without compromising expert selectivity.
-Extensive evaluations conducted on seven benchmark datasets—MedQA, CoLA, SST-2, GSM8K, ARC-E, ARC-C, and HumanEval—demonstrate that integrating LoRA-Mixer substantially improves model performance across all evaluated tasks.
-Additionally, cross-domain experiments confirm the versatility and adaptability of our proposed framework.
-Compared to state-
+Extensive evaluation on 15 benchmark datasets (MedQA, GLUE, GSM8K, ARC-E, ARC-C, HumanEval, PIQA, HellaSwag, and BoolQ) demonstrates that integrating LoRA-Mixer significantly improves model performance across all evaluated tasks.
+With only 48% of the parameters of existing methods, the RSL-optimized LoRA-Mixer outperforms the state-of-the-art routing and LoRA-MoE baseline models, achieving improvements of +3.79%, +2.90%, and +3.95% on GSM8K, CoLA, and ARC-C, respectively.
+Cross-model transfer and adaptive module reuse experiments further demonstrate the method’s versatility and data efficiency.
+2 Related Work.
+PEFT For LLMs Low-rank adaptation (LoRA) Hayou et al.
+(2024); Zhou et al.
+(2024); Hu et al.
+(2022); Xu et al.
+(2024); Sheng et al.
+(2023); Biderman et al.
+(2024); Li et al.
+(2023); Zeng et al.
+(2025); Lin et al.
+(2025)
 Figure 2: The overall architecture of LoRA-Mixer.
@@
 The routing training is guided by RSL loss for balancing experts loads and specificity.
-of-the-art approaches, LoRA-Mixer achieves significant efficiency, using only 48% of the training parameters while improving performance by 1.09%, 1.45%, 1.19%, and 0.85% on GSM8K, HumanEval, SST-2, and CoLA benchmarks, respectively.
-2 Related Work
-PEFT For LLMs Low-rank adaptation (LoRA) [12, 11, 4, 10, 47, 49, 50, 51] effectively fine-tunes large models by learning a low-rank matrix and freezing the original weights.
+effectively fine-tunes large models by learning a low-rank matrix and freezing the original weights.
 While effective for a single task, its task-specific nature limits generalization.
-Recent work combines LoRA with mixture of experts (MoE) [13, 14, 25, 15, 16] to achieve dynamic adaptation.
-For example, MixLoRA [13] uses LoRA experts for top-k routing in FFN, improving multi-task performance but suffering from gradient entanglement issues.
-MoLE [14] combines LoRA layers via gating but lacks sparse routing.
-LoraHub [15] performs gradient-free few-shot combination of LoRA modules for unseen tasks but struggles with complex semantics due to lack of gradient optimization and dynamic routing.
-Other methods [18, 19, 45, 46, 48, 60] explore flexible routing mechanisms to improve the model’s adaptability.
+Recent work combines LoRA with mixture of experts (MoE) Li et al.
+(2024a); Wu et al.
+(2024); Ouyang et al.
+(2025); Huang et al.
+(2023); Zhao et al.
+(2024); Song et al.
+(2022; 2023); Hu et al.
+(2025) to achieve dynamic adaptation.
+For example, MixLoRA Li et al.
+(2024a) uses LoRA experts for top-k routing in FFN, improving multi-task performance but suffering from gradient entanglement issues.
+MoLE Wu et al.
+(2024) combines LoRA layers via gating but lacks sparse routing.
+LoraHub Huang et al.
+(2023) performs gradient-free few-shot combination of LoRA modules for unseen tasks but struggles with complex semantics due to lack of gradient optimization and dynamic routing.
+Other methods Li et al.
+(2025); Huang et al.
+(2025); Yang et al.
+(2024); Chen et al.
+(2023); Luo et al.
+(2024); Li et al.
+(2024b); cite11 explore flexible routing mechanisms to improve the model’s adaptability.
 However, these methods usually introduce additional routing networks or optimization targets, resulting in instability during training, limiting their application in actual multi-task or low-resource scenarios.
@@
 By selectively activating a subset of expert modules for each input, MoE allows the model to scale capacity without linearly increasing the amount of computation.
-As a result, more and more large models have adopted MoE, including GLaM[20], Switch Transformers[6], and the recent DeepSeek series[21, 22].
+As a result, more and more large models have adopted MoE, including GLaMDu et al.
+(2022), Switch TransformersFedus et al.
+(2022); Song et al.
+(2025); Zhang et al.
+(2025a; b); Ye et al.
+(2025), and the recent DeepSeek seriesLiu et al.
+(2024); Guo et al.
+(2025a).
 These advances indicate that MoE is becoming a mainstream architectural trend in the development of next-generation base models.
-Among them, GLaM[20] and Switch Transformer[6] build a mixture of experts (MoE) model in the FFN module, and use a sparse activation mixed expert architecture to expand the model capacity and achieve better performance.
-LLaVA-MoLE[23] uses a top-1 strategy to route tokens to domain-specific expert models, thereby alleviating data conflicts and achieving continuous performance improvements over the ordinary LoRA baseline.
-LoRAMoE[24] uses routers to integrate LoRA experts while retaining general knowledge.
-HMoRA[35] combines the layered fine-tuning methods of MoE and LoRA, and gradually switches the routing strategy as the number of layers increases.
-MoLA[34] assigns different numbers of experts at different levels, proving that deeper layers often require more experts.
-In our method, we use mixed LoRA experts in the projection layer and optimize the load loss.
-LoRA-Mixer is a more fine-grained MoE construction method that is independent of the architecture and can simultaneously maintain expert expertise, computational sparsity, and routing adaptability.
-In addition, LoRA-Mixer can reuse existing LoRA modules and only requires very little data to train routing.
-While saving computing resources, it achieves the expansion of model capacity and the improvement of generalization ability.
-3 Method
-In this section, we introduce LoRA-Mixer, a flexible and pluggable mixture of experts (MoE) framework for combing multiple LoRA experts of LLMs.
+Among them, GLaMDu et al.
+(2022) and Switch TransformerFedus et al.
+(2022) build a mixture of experts (MoE) model in the FFN module, and use a sparse activation mixed expert architecture to expand the model capacity and achieve better performance.
+LLaVA-MoLEChen et al.
+(2024) uses a top-1 strategy to route tokens to domain-specific expert models, thereby alleviating data conflicts and achieving continuous performance improvements over the ordinary LoRA baseline.
+LoRAMoEDou et al.
+(2023) uses routers to integrate LoRA experts while retaining general knowledge.
+HMoRALiao et al.
+(2025) combines the layered fine-tuning methods of MoE and LoRA, and gradually switches the routing strategy as the number of layers increases.
+MoLAGao et al.
+(2024) assigns different numbers of experts at different levels, proving that deeper layers often require more experts.
+Notably, we only utilize hybrid LoRA experts in the core projection layer, resulting in extremely high parameter efficiency.
+By applying RSL to LoRA-Mixer, we can efficiently reuse LoRA modules and require minimal training data, saving computational resources while expanding model capacity and improving generalization.
+3 Method.
+In this section, we introduce LoRA-Mixer, a flexible and pluggable Mixture of Experts (MoE) framework for combining LoRA experts from multiple LLMs.
 3.1 Preliminaries
-LoRA [4] is a parameter-efficient fine-tuning method that adapts large pre-trained models by introducing low-rank updates to the original weight matrices, instead of updating them directly.
-Given a pre-trained weight matrix $W_{0}\in\mathbb{R}^{d_{\text{out}}\times d_{\text{in}}}$, LoRA freezes $W_{0}$ and introduces a trainable low-rank decomposition $\Delta W=BA$, where $A\in\mathbb{R}^{r\times d_{\text{in}}}\,,\ B\in\mathbb{R}^{d_{\text{out}}\times r}\,,\ r\ll\min(d_{\text{in}},d_{\text{out}})$.
-The adapted transformation becomes:
-$y=(W_{0}+BA)\mathbf{x}=W_{0}x+B(A\mathbf{x}),$ (1)
-where $\mathbf{x}\in\mathbb{R}^{d_{\text{in}}}$ is the input token representation.
-This technique significantly reduces trainable parameters to $r(d_{\text{in}}+d_{\text{out}})$, enabling scalable fine-tuning with limited resources.
-Mixture-of-Experts [3] is a sparse neural architecture where each input token is processed by a small subset of expert networks.
+Mixture-of-Experts is a sparse neural architecture where each input token is processed by a small subset of expert networks.
 Given $K$ experts and a router that produces a score vector $G(\mathbf{x})\in\mathbb{R}^{K}$, a softmax is applied to obtain the routing distribution:
@@
 This design allows MoE to reduce computational cost while enabling experts to specialize on different input patterns.
+Auxiliary loss. To encourage balanced expert utilization, MoE training introduces an auxiliary loss. We define the expected routing probability and top-1 usage as
+$\bar{p}_{i}=\mathbb{E}_{x\sim\mathcal{D}}[p_{i}(x)],\qquad\bar{f}_{i}=\mathbb{E}_{x\sim\mathcal{D}}\big[\mathbb{I}(i=\arg\max_{j}p_{j}(x))\big],\qquad L_{\text{aux}}=\alpha\sum_{i=1}^{K}\bar{p}_{i}\,\bar{f}_{i}.$ (3)
+where $\bar{p}_{i}$ denotes the average routing intention and $\bar{f}_{i}$ the empirical usage. We demonstrate in Appendix A.17 why the auxiliary loss leads to over-averaging.
 3.2 LoRA-Mixer for Compositing LoRAs
@@
 Our proposed LoRA-Mixer framework is highly flexible and supports the integration of LoRA modules from diverse sources.
-In one common scenario, users may download pre-trained LoRA adapters from public repositories such as LoRAHub [15], which currently hosts 196 high-quality LoRA modules across a wide range of domains.
+In one common scenario, users may download pre-trained LoRA adapters from public repositories such as LoRAHub Huang et al.
+(2023), which currently hosts 196 high-quality LoRA modules across a wide range of domains.
 These can be directly composed using LoRA-Mixer with minimal additional data.
@@
 3.3 Specialization Balance Loss for Routing Optimization
-The next step is to optimize the expert router.
-Although previous studies  [22, 21, 13] introduced auxiliary losses to align the average gating score with the expert utilization to promote load balancing, we observed that this approach overemphasized consistency, resulting in an overly balanced distribution of experts.
-In this case, all experts are forced to be used equally regardless of the input semantics.
-This hinders effective routing and often requires more training data.
-To ensure that the expert load is balanced and the routing is input-aware, we propose an improved optimization objective called Route-Specialization Balance Loss (RSL).
-We introduce a selectivity-aware regularization term that regulates the entropy of routing distribution to enhance the auxiliary loss, guiding routers to make more discriminative expert choices instead of blindly averaging activations.
-Formally, let $\bar{p}_{i}$ denote the average soft route score (across tokens) and $\bar{f}_{i}$ denote the normalized score of the token assigned to expert $i$ in the first $k$ routes.
+The next step is to optimize the expert routers.
+While previous research has introduced auxiliary loss functions to align the average gating score with expert utilization, thereby promoting load balancing, we observed that this approach overemphasizes consistency, resulting in an overly balanced expert distribution.
+In this case, all experts are forced to be used equally, regardless of the input semantics.
+This hinders effective routing and generally requires more training data.
+We propose a new perspective: viewing the router as an information bottleneck that determines the degree to which semantic differences between tokens are preserved or compressed.
+From this perspective, the entropy of the routing distribution quantifies the uncertainty of the router.
+Therefore, load balancing and specialization selection appear to be in conflict.
+To achieve a balance between the two, ensuring balanced expert load and input-aware routing, we propose an improved optimization objective, called Routing Specialization Balance Loss (RSL).
+RSL addresses this issue by unifying the global budget and local selectivity into a single objective.
+Specifically, RSL enhances the auxiliary loss with an entropy regularization term, improving the accuracy of expert selection by suppressing overly flat distributions.
+Formally, let $\bar{p}_{i}$ denote the average soft routing score (across tokens) and $\bar{f}_{i}$ denote the normalized score assigned to the token of expert $i$ in the first $k$ routes.
 The RSL loss function is defined as:
 $\mathcal{L}_{\text{RSL}}=\alpha\cdot\sum_{i=1}^{K}\bar{p}_{i}\cdot\bar{f}_{i}-\lambda\cdot\mathbb{E}_{x\sim\mathcal{D}}\left[\mathcal{H}(p(\mathbf{x}))\right],$ (5)
-Where $\alpha$ controls the strength of the balanced consistency term, $\lambda$ is a small positive coefficient of the entropy regularizer:
+Where $\alpha$ controls the strength of the equilibrium consistency term, $\lambda$ is a small positive coefficient of the entropy regularization term:
 $\mathcal{H}(p(\mathbf{x}))=-\sum_{i}p_{i}(\mathbf{x})\log p_{i}(\mathbf{x})$ (6)
-which is calculated based on the routing distribution $p(\mathbf{x})$ of each token.
-The first term enforces that routing intentions are consistent with actual usage, while the second term penalizes excessively high entropy in expert selection, thereby promoting more specialized expert assignments.
-This combination ensures that the router not only maintains overall balance, but also forms a preference structure that avoids the problem that every expert is used equally regardless of the input semantics.
-We have analyzed RSL and traditional auxiliary loss from a theoretical perspective.
-Please refer to Appendix C for detailed derivation.
-For the balance loss in training, please refer to the Appendix  D.
+The design principles behind RSL are: 1) From an information bottleneck perspective, minimizing $\mathcal{H}(p(x))$ reduces token-conditional uncertainty under a fixed global load, directly promoting specialization without disrupting the balance.
+2) The entropy term acts as a curvature provider in the routing simplex, resulting in strong convexity and well-conditioned optimization (see A.1); this is precisely where traditional auxiliary loss functions are weak.
+3) The coefficient $\lambda$ becomes an interpretable knob that trades off global fairness and local specialization, rather than simply a regularization weight.
+Specifically, the entropy term introduces token-specific entropy. We derive the entropy gradient with respect to the routing score $p_{i}(x)$:
+$\frac{\partial\mathcal{H}(p(x))}{\partial p_{i}(x)}$ $=-\log p_{i}(x)-1,\quad\text{subject to}\quad\sum_{i=1}^{K}p_{i}(x)=1,$ (7)
+$=-\log p_{i}(x)-1+\mu,$ (8)
+where $\mu$ is the Lagrange multiplier due to the simplex constraint. Therefore, the total gradient of the RSL loss becomes:
+$\nabla_{p_{i}(x)}\mathcal{L}_{\text{RSL}}=\alpha\cdot\frac{\partial\bar{p}_{i}}{\partial p_{i}(x)}\cdot\bar{f}_{i}+\lambda(\log p_{i}(x)+1-\mu).$ (9)
+This demonstrates that RSL introduces a token-level gradient signal via $\log p_{i}(x)$, unlike the auxiliary loss, which only propagates global gradients.
+Specifically, the $\log p_{i}(x)$ term provides a token-level signal that amplifies the input-aware variance in the expert’s choices, thereby counteracting the bias towards consistency in global balance.
+To quantify input-aware routing, we define:
+$\text{Var}_{x\sim\mathcal{D}}(p(x)):=\mathbb{E}_{x}\left[\|p(x)-\bar{p}\|^{2}\right].$ (10)
+We call routing token-aware if $\text{Var}(p(x))>\epsilon$ and $\epsilon>0$.
+The auxiliary loss function tends to reduce this variance (driving uniform routing), while RSL encourages high variance and peaked distributions that are consistent with the input semantics.
+In addition to the above proofs, we provide convergence analysis and generalization bound proofs in the Appendix A.1 A.2.
+The former introduces strong convexity on the product simplex and proves that the added entropy term can maintain stable optimization dynamics; the latter shows that entropy regularization reduces the assumed complexity of the router and improves robustness in low-data environments.
 Routing Optimization.
-After we prepare all LoRAs, we apply a soft training process on router.
-To prevent the expert knowledge in the first phase from being contaminated, we introduce a regularization term to penalize deviations from the previously learned expert parameters.
-Let the first phase parameters of expert $i$ be $\theta_{i}^{(0)}$ and the current parameters be $\theta_{i}$.
+After LoRA is ready, we soft-train the router.
+To prevent expert knowledge from being contaminated, we introduce a regularization term to penalize deviations from the previously learned expert parameters.
+Let the first-stage parameters of expert $i$ be $\theta_{i}^{(0)}$ and the current parameters be $\theta_{i}$.
 We define the regularization term as:
@@
 where $\mathcal{C}$ represents the set of constrained experts and $\beta$ controls the regularization strength.
-This regularization term constrains the sensitive experts to stay close to their original knowledge while still allowing other experts to adjust.
-For complex tasks, we support expert-level control, thus enabling flexible multi-expert learning.
-To ensure that all experts can obtain meaningful gradients during the joint training process and promote stable optimization of the routing balance loss, we adopt soft expert fusion in training.
-Specifically, the router outputs the softmax routing scores $\mathbf{p}_{b,t}\in\mathbb{R}^{K}$ of all experts and fuses them, thereby achieving a fully differentiable mixture of LoRA experts.
-Although soft routing provides stable optimization and gradient propagation for all experts, its combination with the auxiliary loss leads to the problem of extremely balanced expert usage, that is, all experts are activated to the same extent regardless of the input semantics.
-To address this limitation, we introduce the Route-Specialization Balancing Loss (RSL) 3.3, which promotes the specialization of experts without sacrificing load balancing.
-The total loss during joint training becomes:
-$\mathcal{L}_{\text{total}}=\underbrace{\mathcal{L}_{\text{task}}}_{\text{Supervised Objective}}+\underbrace{\alpha\cdot\mathcal{L}_{\text{RSL}}}_{\text{Routing Specialization}}+\underbrace{\beta\cdot\sum_{i\in\mathcal{C}}\left\|\theta_{i}-\theta_{i}^{(0)}\right\|^{2}}_{\text{Expert Preservation}},$ (8)
+This regularization term constrains sensitive experts to maintain their original knowledge while allowing other experts to flexibly adjust, supporting multi-expert learning for complex tasks.
+To ensure effective gradients and stable optimization during joint training, we employ soft expert fusion: the routers output a softmax score $\mathbf{p}_{b,t}\in\mathbb{R}^{K}$, achieving differentiable LoRA hybridization.
+While this approach provides stable gradients, combining it with an auxiliary loss results in equal activations among experts, inhibiting specialization.
+To address this, we introduce RSL, which promotes expert differentiation while maintaining load balancing and enables efficient reuse of LoRA with minimal data.
+The total loss during joint training is:
+$\mathcal{L}_{\text{total}}=\mathcal{L}_{\text{task}}+\alpha\cdot\mathcal{L}_{\text{RSL}}+\beta\cdot\sum_{i\in\mathcal{C}}\left\|\theta_{i}-\theta_{i}^{(0)}\right\|^{2},$ (12)
 where $\mathcal{L}_{\text{task}}$ denotes the standard task loss (e.g., cross-entropy loss).
-$\alpha$ is weighting factor for RSL.
+$\mathcal{L}_{\text{RSL}}$ is the routing specialization balance loss, which is used to promote consistency between routing scores and actual token assignments.
+$\alpha$ is a weighting factor for RSL.
 $\mathcal{L}_{\text{preserve}}$ is the expert regularization loss described above, scaled by $\beta$.
-At inference time, we adopt a Top-3 routing strategy.
-By separating soft fusion at training time from sparse activation at inference time, we can ensure that the model obtains robust gradient propagation during routing optimization while performing efficient inference.
-4 Experiment
-To evaluate the effectiveness of our proposed LoRA-Mixer framework, we conduct experiments with LoRAs finetuned in five domains: Medical QA, Commonsense Reasoning, Natural Language Understanding, Mathematical Reasoning, and Coding Ability.
+We provide an exploration of how to choose hyperparameters in Appendix A.8.
+4 Experiment.
+In this section, we conduct extensive experiments on 15 datasets across five domains (MedQA, commonsense reasoning, NLP, mathematics, and coding).
 4.1 Experimental Setup
-Datasets.
-To evaluate LoRA-Mixer, we selected seven publicly available benchmarks.
-For medical question answering, we used the Medical-QA dataset.
-For commonsense reasoning, we adopted ARC-E [30] and ARC-C [30], both of which focus on multiple-choice questions.
-For natural language understanding, we used SST2 [31] and CoLA [31], where SST2 is used for sentiment classification and CoLA is used for grammatical judgment.
-For mathematical reasoning, we chose the GSM8K [29] dataset, which contains thousands of elementary school math problems that require multi-step solutions.
-Finally, to evaluate encoding ability, we chose the HumanEva [19]l dataset.
-Baselines.
-We chose three open source LLMs - LLaMA3-8B [26], Mistral-7B [27] and Falcon-Mamba-7B [28].
-LLaMA3-8B and Mistral-7B both rely on the Transformer backbone network; Falcon-Mamba-7B is built entirely on the Mamba paradigm.
-In addition, we compared LoRA-Mixer with other state-of-the-art methods, including MoLE [14],MixLoRA [13],LoraHub [15],LoRA-LEGO [16].
-Evaluation metrics We use three evaluation metrics to measure the performance.
-Specifically, for GSM8K [29], ARC-E [30], ARC-C [30], SST2 [31], and CoLA [31], we use ACC to measure performance.
-For the HumanEval dataset [19], we adopt the $Pass@1$ metric, which represents the ability of a single generated answer to correctly solve the task.
-Finally, considering the domain-specific freedom and rigor required by the Medical-QA dataset, we use DeepSeek-R1 [22] to evaluate completeness, correctness, and logical clarity, and report the final percentage scores.
+Datasets We tested RSL and LoRA-Mixer on 15 public benchmarks: MedicalQA, ARC-E, ARC-C, GSM8K, GLUE, PIQA, BoolQ, HumanEval, and HellaSwag.
+These datasets cover five different domains.
+For detailed dataset information, please refer Appendix 11.
+Baseline We selected LLaMA3-8B, Mistral-7B, and Falcon-Mamba-7B as the basemodel.
+LLaMA3-8B and Mistral-7B are Transformer architectures, while Falcon-Mamba is a pure SSM architecture.
+We compare LoRA-Mixer and RSL with other state-of-the-art methods, including MoLE Wu et al.
+(2024), MixLoRA Li et al.
+(2024a), LoraHub Huang et al.
+(2023), LoRA-LEGO Zhao et al.
+(2024) and PHATGOOSE Muqeeth et al.
+(2024).
+We also select strong baselines that specifically optimize auxiliary losses to demonstrate the effectiveness of RSL, including GMoE Bai et al.
+(2024), AESL Guo et al.
+(2025b), and DsMoE Pan et al.
+(2024).
+It is worth noting that MoLE and MixLoRA also optimize auxiliary losses.
+For parameter, training and inference analysis, please refer to A.4 A.7.
+Metrics For HumanEval, we use the $Pass@1$ metric.
+Considering the domain-specific freedom and rigor required by the Medical-QA dataset, we use DeepSeek-R1 for evaluation.
+For the remaining tasks, we use ACC as the metric.
+To reduce randomness, all experiments are run three times and the average reported.
 4.2 Comparisons
-Table1 shows the performance indicators of the three basemodels on various tasks. To ensure the reliability of the experimental results, we ran all experiments three times and took the average value.
 Table 1: Performance of three base models, Falcon-Mamba-7B, Mistral-7B, and LLaMA3-8B, on seven benchmarks.
 Base Model Medical CoLA SST2 GSM8K ARC-E ARC-C HumanEval
-Falcon-Mamba-7B[28] 73.67 82.42 92.81 52.54 77.61 68.78 29.29
-Mistral-7B[27] 66.32 71.21 85.24 40.83 80.00 61.50 27.95
-LLaMA3-8B[26] 78.47 79.14 93.12 57.92 88.45 78.65 52.44
-We compare our method with the state-of-the-art methods [14, 13, 15, 16] on seven datasets, and the experimental results are shown in Table 2.
-As can be seen from Table 2, our method achieves better performance than the baseline on most tasks.
-For the Falcon-Mamba , our method significantly outperforms other baselines on all tasks.
-For model details, please refer to Appendix B.
-Table 2: Comparison of our LoRA-Mixer with LoRAHub [15], MoLE [14], and Mix-LoRA [13] across seven tasks (best scores in bold).
-Note that MixLoRA is excluded from Falcon-Mamba due to its Transformer-specific design.
+Falcon-Mamba-7B 73.67 82.42 92.81 52.54 77.61 68.78 29.29
+Mistral-7B 66.32 71.21 85.24 40.83 80.00 61.50 27.95
+LLaMA3-8B 78.47 79.14 93.12 57.92 88.45 78.65 52.44
+Table1 shows the performance indicators of the three basemodels on various tasks.
+We compare our approach with the state-of-the-art methods on seven datasets.
+The experimental results are shown in Table 2.
+Our approach outperforms the baselines on most tasks.
+For the Falcon-Mamba dataset, our approach significantly outperforms the baselines on all tasks.
+For model details, see Appendix A.11.
+We use the LLaMA2-7B from the LoRA-LEGO paper as the basemodel and conduct experiments on the four tasks: CoLA, SST2, MRPC, and RTE.
+The LoRA configuration uses $r=6$ and $\alpha=12$.
+The experimental results are shown in Table 4.
+As can be seen, our method outperforms LoRA-LEGO on three of the four tasks.
+Table 2: Comparison of our LoRA-Mixer with LoRAHub, MoLE, and Mix-LoRA across seven tasks (best scores in bold). Note that MixLoRA is excluded from Falcon-Mamba due to its Transformer-specific design.
 Metho Medical CoLA SST2 GSM8K ARC-E ARC-C HumanEval
 Falcon-Mamba (7B)
-LoRAHub [15] 70.14 81.11 93.35 51.64 81.16 72.37 30.68
-MOLE [14] 74.51 84.77 94.22 54.28 83.46 76.61 33.57
+LoRAHub 70.14 81.11 93.35 51.64 81.16 72.37 30.68
+MOLE 74.51 84.77 94.22 54.28 83.46 76.61 33.57
 LoRA 77.26 85.62 95.07 56.27 85.68 76.51 33.54
@@
 Mistral (7B)
-LoRAHub [15] 69.17 75.73 90.21 44.94 81.14 69.21 32.60
-MOLE [14] 71.07 78.51 94.17 45.31 85.68 68.77 35.37
-MixLoRA [13] 69.74 78.61 93.44 45.50 85.42 69.15 33.80
+LoRAHub 69.17 75.73 90.21 44.94 81.14 69.21 32.60
+MOLE 71.07 78.51 94.17 45.31 85.68 68.77 35.37
+MixLoRA 69.74 78.61 93.44 45.50 85.42 69.15 33.80
 LoRA 70.33 79.19 93.58 46.67 86.66 70.53 35.31
@@
 LLaMA-3 (8B)
-LoRAHub [15] 78.11 79.84 92.77 59.10 87.13 80.14 52.83
-MOLE [14] 78.43 81.37 94.18 63.81 88.15 81.77 55.87
-MixLoRA [13] 79.87 80.67 94.22 64.44 88.70 82.90 55.49
+LoRAHub 78.11 79.84 92.77 59.10 87.13 80.14 52.83
+MOLE 78.43 81.37 94.18 63.81 88.15 81.77 55.87
+MixLoRA 79.87 80.67 94.22 64.44 88.70 82.90 55.49
 LoRA 81.09 81.50 95.30 65.14 89.59 82.15 55.61
 LoRA-Mixer(ours) 81.55 82.22 95.41 65.53 89.88 83.24 57.32
-Table 3: Evaluation of LoRA-Mixer on LoRAs sourced from Internet on five GLUE tasks, the base model is Flan-T5[36].
+Table 3: Evaluation of LoRA-Mixer on LoRAs sourced from Internet on five GLUE tasks, the base model is Flan-T5Chung et al. (2024).
 Method SST-2 CoLA MRPC RTE QQP
-Flan-T5 [36] 94.01 74.21 79.90 80.08 82.32
+Flan-T5 94.01 74.21 79.90 80.08 82.32
 LoRA 94.50 80.54 83.76 83.47 85.55
 LoRA-Mixer 95.07 82.14 85.15 85.31 84.75
-Table 4: Comparison of LoRA-Mixer and LoRA-LEGO. Results for LoRA-LEGO are from its paper [16].
+Table 4: Comparison of LoRA-Mixer and LoRA-LEGO. Results for LoRA-LEGO are from its paper.
 Method CoLA SST-2 MRPC RTE
 LoRA 61.63 75.74 68.00 52.22
-LEGO [16] 55.48 73.22 66.00 71.85
+LEGO 55.48 73.22 66.00 71.85
 LoRA-Mixer 64.60 80.31 72.24 61.47
-Since LoRA-LEGO [16] is not open-sourced, to compare fairly, we use LLaMA2-7B [33] as the base model and conduct experiments on four tasks including CoLA, SST2, MRPC and RTE from LoRA-LEGO paper.
-Lora’s configuration employs a low rank of $r=6$ and a scaling factor of $\alpha=12$.
-The experimental results are shown in Table 4.
-From the results, it can be seen that our method outperforms LoRA-LEGO in three of the four tasks.
-Considering that Mistral-7B and LLaMA3-8B have exactly the same model architecture, we directly migrate the parameters trained on Mistral-7B to LLaMA3-8B without any fine-tuning and adaptation, and conduct experiments on three datasets: ARC-E, ARC-C, and GSM8K.
+Considering that Mistral-7B and LLaMA3-8B have the same architecture, we directly migrate the parameters trained on Mistral-7B to LLaMA3-8B without any fine-tuning and adaptation, and conduct experiments on three datasets: ARC-E, ARC-C, and GSM8K.
 The results are shown in Table 5 .
@@
 + Mistral 59.13 (1.02) 76.26 (1.01) 81.43 (1.04) 85.89 (0.97) 79.14 (1.01)
-Interestingly, we observed that we achieved better performance than the LLaMA3-8B baseline on two of the three tasks.
-This cross-model migration verifies that the LoRA expert and the learned routing function are not tightly coupled with a specific base model, making it possible to share experts between models with the same architecture.
+Interestingly, we observe that we outperform the LLaMA3-8B on two of the three tasks.
+This cross-model transfer validates the design motivation of LoRA-Mixer and demonstrates that the routing learned via RSL is extremely robust and transferable, making it possible to share experts between models with the same architecture.
+To further demonstrate the performance of LoRA-Mixer, we conducted experiments on BoolQ, HellaSwag, and PIQA using LLaMA3-8B, with $r=32$ for LoRA.
+The results are shown in Table 7.
+It can be seen that LoRA-Mixer can effectively coordinate the capabilities of LoRAs to enhance the basemodel.
+To test the generalization of LoRA-Mixer, we conducted experiments with PHATGOOSE on 9 datasets, six of which were in-distribution experiments and three were out-of-distribution experiments.
+Here we only show the OOD experimental results because OOD tasks can reflect the generalization ability of LoRA-Mixer.
+For the results of the in-distribution experiment, please refer to A.14.
+The results show that LoRA-Mixer has excellent generalization ability, which means that through RSL, routing can make professional decisions and achieve better performance.
+Table 6: OOD comparison across datasets for Phatgoose, and LoRA-Mixer.
+Dataset Base Phatgoose LoRA-Mixer
+QQP 68.84 69.28 69.47
+RTE 70.88 69.87 71.31
+MRPC 74.73 74.59 74.93
+Table 7: Results on BoolQ, HellaSwag, and PIQA. Base is LLaMA3-8B.
+Dataset Base LoRA LoRA-Mixer Gap
+BoolQ 71.25 74.46 79.37 +4.91
+HellaSwag 75.33 77.39 82.41 +5.02
+PIQA 78.47 80.71 84.94 +4.23
 4.3 Testing on LoRAs Sourced from Internet
-To verify flexibility and plugin and play nature of LoRA-Mixer, we test our LoRA-Mixer on LoRAs sourced from Internet.
-Specifically, we download five distinct LoRAs from LoRAHub [15], which were trained on SST2, CoLA, MRPC, RTE, and QQP, respectively (please refer to Appendix E for more details).
-We use the Flan-T5 [36] model as a base model and mount the LoRAs without any modification, and collect $2K$ mixed data for routing training.
+To verify that LoRA-Mixer can reuse LoRAs with minimal effort, we tested it on LoRAs from the internet.
+We downloaded five different LoRAs from LoRAHub Huang et al.
+(2023) and trained them on SST2, CoLA, MRPC, RTE, and QQP (see Appendix A.15 for details).
+We used the Flan-T5 model as the base model, froze the LoRAs parameters, and collected an additional 2k mixed data points for routing training, which is independent of the LoRAs training data.
 The results are shown in Table 4.
-LoRA-Mixer achieved better performance on the four tasks, confirming that LoRA-Mixer’s potential for product-ready multitask applications.
+LoRA-Mixer achieved excellent performance on all four tasks, demonstrating its potential for production-grade multi-task applications.
+4.4 Comparison with other optimized routing losses.
+Although we have compared RSL with other optimized routing losses (MoLE, MixLoRA), to further demonstrate the effectiveness of RSL, we choose three strong baselines for comparison: GMoE, Ds-MoE, and AESL, all of which are specifically optimized for routing losses.
+It is worth noting that all experiments are conducted with the same training data (2k), and the only difference is the routing loss.
+As shown in Table 4.4, RSL significantly outperforms other strong baselines in low-data-resource scenarios.
+Table 8: Comparison with GMoE, DS-MoE, and AESL under the same training data (2k) and LoRAs parameters. Demonstrates the advantages of RSL under low data resources.
+Task GMoE DS-MoE AESL RSL
+SST-2 91.38 92.45 92.64 95.41
+CoLA 79.57 79.83 80.42 82.22
+ARC-E 85.65 85.32 86.24 89.88
+ARC-C 76.42 78.45 79.88 83.24
+HumanEval 46.37 48.92 50.46 57.32
+Figure 3: Expert Assignment Overview.
 4.5 Ablation Study
 Impact of LoRA Rank.
-To evaluate the impact of rank in low-rank adaptation, we conducted experiments on $r=16$, $r=32$, $r=64$, and $r=128$, while keeping all other hyperparameters (e.g., dropout rate, learning rate) unchanged.
+To evaluate the impact of $r$ in LoRAs, we conducted experiments on $r=16$, $r=32$, $r=64$, and $r=128$, while keeping all other hyperparameters (e.g., dropout rate, learning rate) unchanged.
 The results of $r=64$ can be found in Table 2.
-We place the remaining results in Appendix A.
+We place the remaining results in Appendix A.12.
 Expert Load Analysis.
-To analyze the overall load of experts, we uniformly sampled 1K data for seven benchmarks, including Medical, CoLA, SST2, GSM8K, ARC-E, ARC-C, and Humaneval.
-We report the average load of each expert on these $1K$ data, as shown in Figure 4.
-The activation rates of different experts are quite balanced, ranging from 15% - 18%, but in different tasks, the expert load reflects a kind of "perception" ability, and the expert load of specific tasks is higher than that of other experts, as shown in Figure  5.
-This shows that our routing mechanism effectively avoids expert collapse and achieves a balance of expert utilization between different tasks.
-The impact of Top-$K$. To explore the effect of K values on Top-K routing, we used Falcon-Mamba as the basemodel to experiment on SST-2 and CoLA. The experimental results are shown in Figure  3.
-Figure 3: Top-K Routing Impact.
-Figure 4: Expert Assignment Overview.
-As $K$ increases from 1 to 3, we observe that the accuracy of both tasks improves, indicating that using multiple experts allows the model to obtain complementary information.
-However, further increasing the value of K does not guarantee better results, but may degrade the performance.
-Therefore, the setting of $K$ is crucial for the MoE model.
-How to set or dynamically learn the most appropriate $K$ value is a direction worthy of further research in the future.
-Table 6: The impact of LoRA-Mixer optimization on individual LoRAs (LoRA means adding independently trained LoRA, Expert means LoRA optimized by LoRA-Mixer).
-Task LoRA w/o Expert w/ Expert Gap (Expert)
-SST-2 95.30 94.70 95.41 +0.71
-CoLA 81.50 80.15 82.22 +2.07
-GSM8K 65.14 60.17 65.53 +5.36
-HumanEval 55.61 53.39 57.32 +3.93
-Enhanced Expert Performance Analysis.
-In order to verify the enhanced performances of individual expert after LoRA-Mixer optimization.
-We selected four tasks, GSM8K, SST2, CoLA and HumanEval, for experiments.
-The results are shown in Table 6.
-We can find that LoRAs optimized by LoRA-mixer exhibit improved performance, especially in the GSM8K task, where the performance improved by 5.36% after adding mathematical experts.
-This result confirms the enhanced ability of each individual LoRA expert after LoRA-Mixer optimization.
-Cross-domain QA To evaluate the cross-domain generalization ability of LoRA-Mixer, we constructed two question-answering datasets: Medical-Mathematics and Mathematics-Coding.
-Each dataset contains 200 samples generated by DeepSeek-R1.
-These questions are more challenging.
-In the Medical-Mathematics dataset, the model needs to provide effective medical advice and corresponding calculations.
-In the Mathematics-Coding dataset, the model needs to generate correct Python code based on mathematical problems.
-The evaluation results of DeepSeek-R1 are shown in Table  7.
-Table 7: Cross-domain performance of LoRA-Mixer on LLaMA3-8B [26].
-Task Base LoRAHub [15] MOLE [14] MixLoRA [13] LoRA-Mixer
-Math-Medical 69.88 70.53 72.11 72.74 73.41
-Math-Coding 59.37 61.08 62.24 63.10 63.46
-The impact of the RSL. We study the impact of our proposed RSL loss function on the LoRA-Mixer framework. RSL has two major advantages.
-First, it enables routers to achieve global load balancing while maintaining strong input perception, which is a key factor to fully exploit the potential of sparse expert models.
-Second, compared with traditional load balancing loss functions, RSL significantly reduces the amount of training data required for effective router optimization.
-To verify the first conclusion, we conducted experiments on three tasks: Medical, GSM8K, and HumanEval.
-As shown in Figure 5, after using RSL loss, LoRA-Mixer always assigns higher activation weights to experts related to the target task, which reflects the router’s strong domain perception and adaptive specialization capabilities.
-In contrast, using only auxiliary loss, the router will evenly distribute experts regardless of the semantics of the input content, which will result in the potential of the relevant experts not being fully developed, causing a performance bottleneck.
-To support the second conclusion, we analyze the impact of training data size on routing performance.
-Specifically, we construct training sets of different sizes by sampling from a multi-task dataset pool and evaluate the performance on seven benchmark tasks.
-For clarity, we report the average performance over all tasks.
+To analyze the overall load, we uniformly sampled 1K data from seven benchmarks (MedicalQA, CoLA, SST2, GSM8K, ARC-E, ARC-C, and Humaneval).
+We report the average load per expert on this 1K data, as shown in Figure 3 .
+The activation rates of different experts are relatively balanced, ranging from 15% to 18%.
+However, across tasks, expert load reflects a "perception" ability, with experts on certain tasks having higher loads than others, as shown in Figure  4 .
+This demonstrates that our routing mechanism effectively avoids expert collapse and achieves balanced expert utilization across different tasks.
+The impact of K and Cross-domain QA.
+We also analyzed the impact of K in Top-K.
+In general, increasing K within an appropriate range can improve model performance.
+For details, please refer to A.3.
+Furthermore, to evaluate LoRA-Mixer’s generalization capabilities on cross-domain tasks, we created two cross-domain datasets: Mathematics-Coding and Medical-Mathematics.
+For detailed experimental results, please refer to A.3.
+These results demonstrate that LoRA-Mixer maintains strong generalization and input-awareness capabilities on cross-domain datasets.
+The impact of the RSL.
+We investigated the impact of our proposed RSL on the LoRA-Mixer.
+RSL offers three key advantages.
+First, it enables routing to achieve global load balancing while maintaining robust input awareness.
+Second, compared to auxiliary losses, RSL requires less training data.
+Finally, routing trained with RSL is completely independent of the original LoRAs training data.
+To verify the first conclusion, we conducted experiments on Medical, GSM8K, and HumanEval.
+As shown in Figure 4, when using RSL, the router consistently assigns higher activation weights to relevant experts, demonstrating its strong domain awareness and adaptive specialization capabilities.
+In contrast, when using only the auxiliary loss, the router disregards the semantics of the input and evenly distributes experts, resulting in suboptimal performance.
 Table 9: Average performance across seven tasks under different routing training data sizes, with or without RSL. With RSL, LoRA-Mixer requires much less data while showing better performances.
@@
 2K 79.26 77.29 +1.97
-3K 78.64 77.54 +1.10
-4K 78.51 79.28 -0.77
+4K 78.77 79.14 -0.37
+6K 79.41 79.37 -0.04
+8K 79.75 79.48 +0.27
+10K 79.94 79.51 +0.43
 Figure 4: Expert Load Distribution across Tasks.
-As shown in Table 4.4, LoRA-Mixer achieves comparable performance using only 51.62% of the routing supervision data required by traditional auxiliary loss methods.
-We attribute this data efficiency to the dual regularization mechanism introduced by RSL.
-Specifically, a global consistency term $\bar{\mathbf{p}}^{\top}\bar{\mathbf{f}}$ aligns the expected routing probabilities with actual expert utilization, while a local token-level entropy penalty encourages diverse and selective expert activation.
-This synergistic design mitigates the overly uniform expert usage common in auxiliary loss, promoting both expert specialization and routing sparsity.
-As a result, the model maintains robust and adaptive expert assignment even under limited supervision.
-A detailed theoretical justification is provided in Appendix C.
-5 Conclusion and Discussion
-This paper introduces LoRA-Mixer, a flexible and architecture-agnostic MoE framework for combining LoRAs, adapting LLM for multitask.
-It improves the performance of Transformer and SSM models by replacing the projection layer with a dynamically routed LoRA experts.
-Through a two-stage training paradigm, LoRA-Mixer decouples expert learning from routing, enabling specialization and task awareness.
-To address the problem of overly uniform auxiliary losses, we propose RSL, which balances expert load while improving routing selectivity.
-The framework enables efficient router training with minimal data and supports cross-domain reuse of LoRA modules.
+To support the second conclusion, we analyze the impact of training data size on routing performance.
+Specifically, we construct training sets of different sizes by sampling from a multi-task dataset pool and evaluate the performance on seven benchmark tasks.
+For clarity, we report the average performance over all tasks.
+As shown in Table 4.5, RSL achieves comparable or even superior performance using only 51.62% of the training data with auxiliary loss.
+We explain the suboptimal RSL results at 4k in A.16.
+This exceptionally high data efficiency is demonstrated by generalization bound estimation.
+See the Appendix A.2 for a detailed proof.
+5 Conclusion.
+This paper introduces RSL, designed to address the issue of excessively uniform auxiliary loss.
+It achieves highly selective routing while achieving overall load balancing, and enables efficient routing training with minimal data, providing strong support for the reuse of LoRAs modules.
+To verify the effectiveness of RSL, we propose the LoRA-Mixer framework and integrate RSL.
+LoRA-Mixer is a flexible and architecture-agnostic MoE framework for combining LoRAs and adapting LLMs to multitasking.
+Unlike other methods that indiscriminately insert MoEs or completely replace attention or FFN modules, LoRA-Mixer only adapts the core projection layer, improving the performance of Transformers and SSM models.
 Although LoRA-Mixer is effective, its fixed top-$K$ routing may limit adaptability to ambiguous inputs.
@@
 Future work will explore dynamic or differentiable routing and adaptive integration to apply LoRA-Mixer only where most beneficial.
-Appendix A Experiment Result
+6 Acknowledgments
+This research was supported by the National Natural Science Foundation of China (Nos.
+62272184 and 62402189), the China Postdoctoral Science Foundation (Nos.
+2024M751012, 2025T180429, and GZC20230894), and the Hubei Provincial Postdoctoral Research Funding Program (No.
+2024HBBHCXB014).
+The computational work was performed on the high-performance computing platform at Huazhong University of Science and Technology.
+Appendix A Appendix
+A.1 Optimization and Convergence Analysis.
+We analyze the empirical counterpart of $\mathcal{L}_{\text{RSL}}$ on a dataset $S=\{x_{j}\}_{j=1}^{n}$:
+$\widehat{\mathcal{L}}_{\text{RSL}}(S;\{p(x_{j})\})$ $=\;\alpha\sum_{i=1}^{K}\widehat{\bar{p}}_{i}\cdot\widehat{\bar{f}}_{i}\;-\;\frac{\lambda}{n}\sum_{j=1}^{n}\mathcal{H}\!\big(p(x_{j})\big),$
+$\widehat{\bar{p}}_{i}$ $=\frac{1}{n}\sum_{j=1}^{n}p_{i}(x_{j}),\qquad\widehat{\bar{f}}_{i}=\frac{1}{n}\sum_{j=1}^{n}\mathbb{I}\!\big(i=\arg\max_{\ell}p_{\ell}(x_{j})\big).$ (13)
+Smooth surrogate for optimization.
+For optimization analysis we replace the hard top-$1$ frequency by a standard smooth surrogate:
+Assumption 1 (Smoothed usage surrogate).
+Replace $\widehat{\bar{f}}_{i}$ in equation 13 by $\widehat{\bar{s}}_{i}$, where $s_{i}^{\tau}(x)=\mathrm{softmax}(G(x)/\tau)_{i}$ for some $\tau>0$ and $\widehat{\bar{s}}_{i}=\frac{1}{n}\sum_{j}s_{i}^{\tau}(x_{j})$.
+We assume the composite $\sum_{i}\widehat{\bar{p}}_{i}\,\widehat{\bar{s}}_{i}$ is convex and $L$-smooth in $\{p(x_{j})\}$ on the product simplex $\Delta^{n}$.
+[^1] This includes the commonly used choice $\widehat{\bar{s}}_{i}=\widehat{\bar{p}}_{i}$, giving the auxiliary term $\alpha\sum_{i}\widehat{\bar{p}}_{i}^{2}$, which is convex since each $\widehat{\bar{p}}_{i}$ is linear in $\{p(x_{j})\}$.
+Negative-entropy–regularized objective.
+Define
+$F_{S}(\{p(x_{j})\})$ $:=\alpha\sum_{i=1}^{K}\widehat{\bar{p}}_{i}\cdot\widehat{\bar{s}}_{i}+\frac{\lambda}{n}\sum_{j=1}^{n}\sum_{i=1}^{K}p_{i}(x_{j})\log p_{i}(x_{j}).$ (14)
+Since $\sum_{i}p_{i}\log p_{i}=\mathcal{H}(p)$, minimizing $F_{S}$ is identical to minimizing $\widehat{\mathcal{L}}_{\text{RSL}}$ with $\widehat{\bar{f}}_{i}$ replaced by $\widehat{\bar{s}}_{i}$; a change of logarithm base is absorbed into $\lambda$.
+We optimize $F_{S}$ over $\{p(x_{j})\}\in\Delta^{n}$ using entropic mirror descent (exponentiated gradient)
+$p(x_{j})^{t+1}~\propto~p(x_{j})^{t}\odot\exp\!\Big(-\eta_{t}\nabla_{p(x_{j})}F_{S}(\{p(x_{\ell})^{t}\})\Big),\quad\text{then normalize onto }\Delta.$ (15)
+Lemma 1 (Strong convexity).
+The function $h(p)=\sum_{i=1}^{K}p_{i}\log p_{i}$ is $1$-strongly convex on $\Delta$ w.r.t.
+the $\ell_{1}$ norm.
+Hence the regularizer $\frac{\lambda}{n}\sum_{j=1}^{n}h\big(p(x_{j})\big)$ is $\lambda$-strongly convex on the product simplex $\Delta^{n}$.
+Therefore, $F_{S}$ in equation 14 is $\lambda$-strongly convex, since the usage surrogate term is convex by Assumption 1.
+Lemma 2 (Gradient bounds and interior smoothness).
+Under Assumption 1 there exists $G>0$ such that for all feasible $\{p(x_{j})\}$,
+$\Big\lVert\nabla_{p(x_{j})}\Big(\alpha\sum_{i=1}^{K}\widehat{\bar{p}}_{i}\cdot\widehat{\bar{s}}_{i}\Big)\Big\rVert\leq G.$
+Moreover,
+$\Big\lVert\nabla_{p(x_{j})}\frac{\lambda}{n}\sum_{i=1}^{K}p_{i}(x_{j})\log p_{i}(x_{j})\Big\rVert\leq\frac{\lambda}{n}\,\big\|1+\log p(x_{j})\big\|,$
+so if we maintain $p_{i}(x_{j})\geq\epsilon$ during training (e.g., via temperature/clipping), the right-hand side is uniformly bounded by a constant $C(\lambda,\epsilon,K)$.
+On the interior $\{p:p_{i}\geq\epsilon\}$, the entropy term is $L_{\text{ent}}\!\leq\!\frac{\lambda}{n\epsilon}$-smooth (w.r.t.
+$\ell_{2}$), hence $F_{S}$ is $L$-smooth with $L\leq L_{\text{ent}}+L_{\text{aux}}$ for some finite $L_{\text{aux}}=O(\alpha/n)$.
+Theorem 1 (Convergence of EG/EMD for $F_{S}$).
+Under Assumption 1 and Lemmas 12, $F_{S}$ is $\lambda$-strongly convex on $\Delta^{n}$.
+(a) Sublinear rate.
+With steps $\eta_{t}=\frac{1}{\lambda t}$, the averaged iterate $\bar{P}^{\,T}=\frac{1}{T}\sum_{t=1}^{T}P^{t}$ satisfies
+$F_{S}(\bar{P}^{\,T})-F_{S}(P^{\star})\;\leq\;\mathcal{O}\!\Big(\frac{G_{\mathrm{eff}}^{2}}{\lambda\,T}\Big),$
+where $G_{\mathrm{eff}}$ aggregates the uniform gradient bounds in Lemma 2.
+(b) Linear rate on the interior.
+If, in addition, training maintains $p_{i}(x_{j})\geq\epsilon$ (hence $F_{S}$ is $L$-smooth on the interior), then with any constant step $\eta\leq 1/L$ the EG iterates enjoy geometric contraction in KL divergence:
+$D_{\mathrm{KL}}(P^{\star}\!\parallel P^{t+1})\;\leq\;(1-\eta\lambda)\,D_{\mathrm{KL}}(P^{\star}\!\parallel P^{t}),$
+and consequently $F_{S}(P^{t})-F_{S}(P^{\star})$ also decays geometrically.
+The negative-entropy term supplies curvature in the routing simplex (Lemma 1), turning the merely convex surrogate into a strongly convex objective.
+This stabilizes and accelerates routing optimization while preserving input-awareness through the token-level $\log p_{i}(x)$ signals (cf.
+the gradient form in the main RSL, Eq.
+equation 9).
+A.2 Generalization Bound via Algorithmic Stability.
+We analyze generalization under the same smoothed usage surrogate as in optimization (Assumption 1), i.e., replacing $\widehat{\bar{f}}_{i}$ by $\widehat{\bar{s}}_{i}$ (in particular, the commonly used choice $\widehat{\bar{s}}_{i}=\widehat{\bar{p}}_{i}$ yields a convex quadratic auxiliary term).
+We also keep the empirical objective in sample-average form so that the regularization strength does not scale with $n$.
+Let $S=\{x_{j}\}_{j=1}^{n}$ and $S^{(i)}$ be $S$ with the $i$-th example replaced. Denote by $\hat{P}(S)$ a (unique) minimizer of the empirical objective
+$F_{S}(\{p(x_{j})\})~=~\alpha\sum_{i=1}^{K}\widehat{\bar{p}}_{i}\cdot\widehat{\bar{s}}_{i}~+~\frac{\lambda}{n}\sum_{j=1}^{n}\sum_{i=1}^{K}p_{i}(x_{j})\log p_{i}(x_{j}),$
+and similarly $\hat{P}(S^{(i)})$ for $F_{S^{(i)}}$. For reference, the population counterpart is
+$F(\{p(x)\})~=~\alpha\sum_{i=1}^{K}\bar{p}_{i}\cdot\bar{s}_{i}~+~\lambda\,\mathbb{E}_{x}\!\Big[\sum_{i=1}^{K}p_{i}(x)\log p_{i}(x)\Big],\quad\bar{p}_{i}=\mathbb{E}_{x}[p_{i}(x)],~~\bar{s}_{i}=\mathbb{E}_{x}[s_{i}^{\tau}(x)].$
+Assumption 2 (Lipschitz loss & bounded differences).
+(i) The map $\{p(x_{j})\}\mapsto\alpha\sum_{i=1}^{K}\widehat{\bar{p}}_{i}\cdot\widehat{\bar{s}}_{i}$ is $L$-Lipschitz on the product simplex endowed with the averaged block norm $\|P\|_{\text{avg-}1}:=\tfrac{1}{n}\sum_{j}\|p(x_{j})\|_{1}$, and replacing one sample in $S$ changes this auxiliary term by at most $\tfrac{B}{n}$ (bounded-difference property).
+(ii) The per-example evaluation loss $\ell(p;x)$ is $\rho$-Lipschitz in $p$ (under $\|\cdot\|_{1}$ on a single simplex) and bounded in $[0,M]$.
+Lemma 3 (Strong convexity under the averaged product norm).
+Negative entropy $h(p)=\sum_{i=1}^{K}p_{i}\log p_{i}$ is $1$-strongly convex on $\Delta$ w.r.t.
+$\|\cdot\|_{1}$.
+Therefore $\tfrac{\lambda}{n}\sum_{j=1}^{n}h(p(x_{j}))$ is $\lambda$-strongly convex on the product simplex endowed with $\|\cdot\|_{\text{avg-}1}$.
+Since the surrogate usage term is convex (Assumption 1), $F_{S}$ is $\lambda$-strongly convex.
+Lemma 4 (Uniform stability of ERM for $F_{S}$).
+Under Assumptions 1 and 2, let $\hat{P}(S)$ and $\hat{P}(S^{(i)})$ be the ERM solutions. Then the solution mapping is uniformly stable:
+$\big\|\hat{P}(S)-\hat{P}(S^{(i)})\big\|_{\text{avg-}1}~\leq~\mathcal{O}\!\Big(\frac{L+B}{\lambda\,n}\Big).$
+Theorem 2 (Generalization bound for RSL-trained router).
+Under Assumptions 1 and 2, the expected generalization gap satisfies
+$\big|\;\mathbb{E}_{x}[\ell(\hat{P}(S);x)]~-~\tfrac{1}{n}\sum_{j=1}^{n}\ell(\hat{P}(S);x_{j})\;\big|~=~\mathcal{O}\!\Big(\frac{\rho\,(L+B)}{\lambda\,n}\Big),$
+and, with probability at least $1-\delta$,
+$\big|\;\mathbb{E}_{x}[\ell(\hat{P}(S);x)]~-~\tfrac{1}{n}\sum_{j=1}^{n}\ell(\hat{P}(S);x_{j})\;\big|~\leq~\mathcal{O}\!\Big(\frac{\rho\,(L+B)}{\lambda\,n}\Big)~+~\mathcal{O}\!\Big(\sqrt{\tfrac{\log(1/\delta)}{n}}\Big).$
+Lemma 3 gives $\lambda$-strong convexity of $F_{S}$ under $\|\cdot\|_{\text{avg-}1}$.
+By a standard argument for strongly convex ERM with bounded differences, replacing one sample perturbs the objective by at most $O((L+B)/n)$ and yields uniform stability $\beta=\mathcal{O}((L+B)/(\lambda n))$ for the solution map (Lemma 4).
+The Lipschitz evaluation loss then converts stability into generalization: $\mathbb{E}[\text{gap}]\leq\rho\,\beta$, and a standard McDiarmid/Hoeffding argument implies the high-probability term $\mathcal{O}(\sqrt{\log(1/\delta)/n})$.
+This matches the empirical robustness of RSL in low-data regimes: entropy regularization both improves optimization conditioning and reduces the hypothesis sensitivity to single-sample perturbations.
+A.3 Additional experimental results.
+The impact of Top-$K$.
+To explore the impact of Top-K routing, we conducted experiments on SST-2 and CoLA using Falcon-Mamba.
+The results are shown in Figure  5 .
+As $K$ increases from 1 to 3, we observe improved accuracy, suggesting that multiple experts can gain complementary information.
+However, further increasing $K$ may actually degrade performance.
+Therefore, how to set or dynamically learn the optimal $K$ value is a direction worthy of further research.
+Figure 5: Top-K Routing Impact.
+Cross-domain QA To evaluate LoRA-Mixer’s cross-domain generalization capabilities, we constructed two question-answering datasets: Medical-Mathematics and Mathematics-Coding.
+Each dataset contains 200 examples generated by DeepSeek-R1.
+These questions are more challenging.
+We conduct DeepSeek-R1 evaluations, and the results are shown in Table 10.
+Table 10: Cross-domain performance of LoRA-Mixer on LLaMA3-8B.
+Task Base LoRAHub MOLE MixLoRA LoRA-Mixer
+Math-Medical 69.88 70.53 72.11 72.74 73.41
+Math-Coding 59.37 61.08 62.24 63.10 63.46
+A.4 Parameter analysis.
+We tested the trainable parameters of LoRA-Mixer, MoLE, MixLoRA, and LoRAHub under LLaMA3-8B.
+The trainable parameters of LoRA-Mixer and MixLoRA are 3.88% and 8.08%, respectively (six experts + routing).
+LoRA-Mixer’s total parameters account for 48% of MixLoRA’s, with routing parameters accounting for 0.04%.
+MoLE’s training parameters are routing parameters.
+To ensure a fair comparison, we adjusted MoLE’s routing to reduce the trainable parameters to 0.04%.
+LoRAHub is a train-free method with 0% trainable parameters.
+PHATGOOSE’s trainable parameters are gating vectors, specifically 0.01%.
+A.5 Datasets details.
+We present information about all datasets used here for reference.
+Table 11: Statistics of datasets used in experiments.
+Dataset Train Dev Test Task
+ARC-E 2251 570 2376 Multiple-choice QA
+ARC-C 1119 299 1172 Multiple-choice QA
+GSM8K 7473 1319 Math reasoning
+MBPP 779 195 Code generation
+CoLA 8551 1043 1063 Acceptability judgment
+SST-2 10000 872 1821 Sentiment classification
+Medical 10030 4004 4004 Medical QA
+MRPC 3668 408 1725 Paraphrase detection
+MNLI 392702 9815 9832 Natural language inference
+QNLI 104743 5463 5463 Question answering/NLI
+RTE 2490 277 3000 Natural language inference
+QQP 363846 40430 390965 Paraphrase identification
+PIQA 16113 1838 3277 Physical commonsense reasoning
+HellaSwag 39905 10042 10042 Commonsense reasoning
+BoolQ 9427 3270 3270 Boolean question answering
+A.6 Computational Resources and Reproducibility Guarantee.
+Our experiments were conducted on a Linux workstation equipped with a single NVIDIA A800 80GB GPU and a 32-core Intel Xeon CPU, using PyTorch version 2.1.0 and Transformers version 4.36.2.
+For the first stage of training, we used LoRA with parameters r = 64, alpha = 128, dropout = 0.1, batch size = 1, gradient accumulation = 4, and lr = 1e-5.
+The task loss function was cross-entropy loss, and a cosine scheduler with warmup ratio (ratio = 0.1) was used.
+In stage 2, RSL was added with $\alpha=0.1$ and $\lambda=0.5$ in RSL.
+$\beta$ of $\mathcal{L}_{\text{preserve}}$ was fixed to 5e-5, and a top-3 routing strategy was used.
+In the inference phase, we also adopt the top-3 routing strategy and test it in the same device environment.
+A.7 Efficiency and Resource Overhead Analysis.
+We tested inference/training time, and memory usage, and the results are shown in the table 12.
+These metrics highlight the advantages of LoRA-Mixer beyond task performance.
+Inference was performed using a single A800 80G GPU, with max_length=512 and max_new_tokens=15, to test single-sample inference time.
+To eliminate the effects of cold starts, all measurements were performed after a warm-up run.
+Table 12: Inference time comparison across different methods.
+Model Time (s)
+LLaMA3-8B (Baseline) 0.441
+LoRAHub 0.482
+MoLE 0.563
+MixLoRA 0.597
+LoRA-Mixer (Ours) 0.574
+LoRAHub uses a training-free approach, resulting in faster inference speed.
+MoLE is nearly identical to LoRA-Mixer.
+However, this coarse-grained approach often leads to poor performance.
+Overall, LoRA-Mixer achieves a better balance between inference speed and performance.
+Overall, LoRA-Mixer optimized with RSL achieves faster convergence and more stable training.
+For training information, LoRA-Mixer’s training is split into two stages: Stage 1 (expert training) and Stage 2 (router training).
+The results are shown in the table 13.
+Table 13: Training stages with dataset size, time.
+Stage Data Time (hours)
+Stage 1 40k 8.86
+Stage 2 2k 1.74
+Stage 2 4k 3.38
+LoRA-Mixer can skip stage 1 and quickly complete the stage 2 of training using very little data when computing resources are limited, and provide strong performance.
+This is very important in resource-constrained scenarios.
+For memory usage, we measure memory usage for 6 experts, r=64 (A800-80G, float32, batchsize=1, gradientaccumulation=4),the results are shown in the table 14:
+Table 14: Memory usage comparison across different methods.
+Method MixLoRA MoLE LoRA-Mixer
+Memory Usage (GB) 43.07 36.57 38.48
+A.8 Hyperparameter grid search.
+The hyperparameters include $\alpha$$\beta$ and $\lambda$ are not many, and we design a simple yet effective strategy for tuning the hyperparameters.
+We start with heuristic values informed by prior work on multi-objective optimization: $\alpha=0.1$ and $\lambda=0.05$ in $\mathcal{L}_{\text{RSL}}$.
+We fix $\beta=5e-5$ in $\mathcal{L}_{\text{preserve}}$.
+These values are chosen to ensure that the task loss dominates the training in the early stages of training.
+We perform grid search over a narrow range of $\alpha$ and $\lambda$ values on a validation split of all training data (10% of the total data).
+Table 15: Ablation study on $\alpha$ and $\lambda$. Best results are highlighted in bold.
+$\alpha$ $\lambda$ Performance Load Balance (Variance) Routing Entropy
+0.05 0.01 76.2% 0.12 1.9
+0.1 0.05 78.5% 0.08 1.5
+0.2 0.1 75.8% 0.05 1.2
+The values $\alpha=0.1$ and $\lambda=0.05$ achieve the best trade-off across all metrics.
+To prevent RSL from interfering too much with task loss in the early stages of training, we dynamically adjust parameters during training, for example, gradually increasing $\alpha$ and $\lambda$ to their final values in the first 30% of epochs and then keeping them constant.
+This "warm-up" strategy stabilizes training and improves convergence.
+A.9 Additional results on GLUE tasks.
+We present the results of LoRA-Mixer on the QNLI and MNLI tasks, as shown in Table 16.
+The results show that LoRA-Mixer achieves state-of-the-art results on all GLUE tasks, demonstrating the powerful combination of RSL and LoRA-Mixer.
+The experimental settings are LLaMA3-8B and $r=16$, and the average performance over three runs is reported.
+Table 16: Results on QNLI and MNLI. Higher is better.
+Method QNLI MNLI
+Base 90.36 89.95
+LoRAHub 90.44 90.56
+MOLE 91.48 91.04
+MixLoRA 91.73 91.27
+LoRA 91.34 90.87
+LoRA-Mixer (Ours) 92.04 91.45
+We exclude STSB because it focuses on measuring semantic similarity, which is inconsistent with our core goal of "cross-task expert collaboration" (classification, reasoning, and generation), and the results are more likely to be random.
+Furthermore, we want to be consistent with previous studies such as GOAT, LoRA-LEGO, and MixLoRA, which did not conduct experiments on this dataset.
+A.10 Additional Empirical Support for LoRA Migration.
+To better understand LoRA-Mixer’s cross-model transfer experiments, we analyzed the architectures of Mistral-7B and LLaMA3-8B.
+The Mistral-7B and LLaMA3-8B models share a core decoder-specific Transformer structure, and the key dimensions match, as shown below:
+Table 17: Comparison of architecture components between Mistral-7B and LLaMA3-8B.
+Architecture components Mistral-7B LLaMA3-8B
+Number of layers 32 32
+Hidden dimension ($d_{\text{model}}$) 4096 4096
+Attention heads 32 32
+Feedforward dimension ($d_{\text{ff}}$) 14336 14336
+Attention projection layer Q, K, V, output (all $4096\rightarrow 4096$) Q, K, V, output (all $4096\rightarrow 4096$)
+Activation function SwiGLU SwiGLU
+This structural alignment ensures that linear projection layers with the same input/output dimensions can achieve parameter transfer.
+A.11 Falcon-Mamba Architecture Analysis.
+Mamba builds upon the state space model.
+It processes an input sequence $x(t)\in\mathbb{R}^{L}$ to produce an output $y(t)\in\mathbb{R}^{L}$ by employing a hidden state $h(t)\in\mathbb{R}^{N}$.
+This relationship is initially defined by a continuous system:
+$h^{\prime}(t)$ $=\mathbf{A}h(t)+\mathbf{B}x(t),$ (16)
+$y(t)$ $=\mathbf{C}h(t).$
+Here, $\mathbf{A}\in\mathbb{R}^{N\times N}$ is the state transition matrix, and $\mathbf{B}\in\mathbb{R}^{N\times 1}$, $\mathbf{C}\in\mathbb{R}^{N\times 1}$ are projection matrices.
+To process discrete sequences, Mamba discretizes the continuous parameters $\mathbf{A}$ and $\mathbf{B}$ using a time scale parameter $\Delta$ and the zero-order hold (ZOH) principle, resulting in discretized parameters $\mathbf{\overline{A}}$ and $\mathbf{\overline{B}}$:
+$\mathbf{\overline{A}}$ $=\exp\left(\mathbf{\Delta A}\right),$ (17)
+$\mathbf{\overline{B}}$ $=\left(\mathbf{\Delta A}\right)^{-1}\left(\exp\left(\mathbf{\Delta A}\right)-\mathbf{I}\right)\cdot\mathbf{\Delta B}.$
+The discrete state-space equation with a step size of $\Delta$ becomes:
+$h_{t}$ $=\mathbf{\overline{A}}h_{t-1}+\mathbf{\overline{B}}x_{t},$ (18)
+$y_{t}$ $=\mathbf{C}h_{t}.$
+By iteratively expanding the hidden state $h_{t-1}$, Mamba derives a global convolution kernel $\mathbf{\overline{K}}\in\mathbb{R}^{L}$.
+This kernel is then used to compute the output $y$ through a convolution operation with the input $x$:
+$\mathbf{\overline{K}}$ $=\left(\mathbf{C\overline{B},C\overline{A}\overline{B},...,C\overline{A}^{\mathrm{L}-1}\overline{B}}\right),$ (19)
+$y$ $=x\otimes\mathbf{\overline{K}}.$
+Falcon Mamba 7B adopts a pure Mamba architecture, a departure from hybrid designs incorporating staggered attention.
+This deliberate choice aims to maintain the intrinsic linear scalability characteristic of Mamba models.
+To enhance adaptability, the model employs decoupled input embeddings and output weights.
+At its core, Falcon-Mamba features 64 layers of the Falcon-Mamba Mixer.
+Each Mixer layer integrates an SSM (State Space Model) module alongside in-projection and out-projection layers, RMS Norm, and a convolutional layer.
+Within the SSM module, the input is mapped to $\Delta$, $B$, and $C$ through a projection layer denoted as $x$-proj:
+$x\xrightarrow{x\text{-proj}}(\Delta,B,C)$
+where $x$ represents the input to the SSM module. Furthermore, another projection layer, $dt$-proj, discretizes $\Delta$:
+$\Delta\xrightarrow{dt\text{-proj}}\Delta_{discretized}$
+[…]
+A.12 The impact of $r$ in LoRAs
 Table 18: Comparison of LoRA-Mixer on Falcon-Mamba, Mistral, and LLaMA across seven tasks.
@@
 +LoRA-Mixer 80.82 82.25 95.50 63.41 89.33 81.43 56.31
-As shown in Table LABEL:r16,Table 10 and Table  11, within a certain range, as $r$ increases, the performance of the model can be improved to a certain extent.
+As shown in Table LABEL:r16,Table 19 and Table  20, within a certain range, as $r$ increases, the performance of the model can be improved to a certain extent.
 Our method is not only better than the basic model, but even better than the fine-tuned basic model on most tasks.
 This shows that our method can effectively combine existing knowledge through dynamic expert combination to form a more "intelligent" model.
-Appendix B Falcon-Mamba Architecture Analysis
-Mamba builds upon the state space model.
-It processes an input sequence $x(t)\in\mathbb{R}^{L}$ to produce an output $y(t)\in\mathbb{R}^{L}$ by employing a hidden state $h(t)\in\mathbb{R}^{N}$.
-This relationship is initially defined by a continuous system:
-$h^{\prime}(t)$ $=\mathbf{A}h(t)+\mathbf{B}x(t),$ (9)
-$y(t)$ $=\mathbf{C}h(t).$
-Here, $\mathbf{A}\in\mathbb{R}^{N\times N}$ is the state transition matrix, and $\mathbf{B}\in\mathbb{R}^{N\times 1}$, $\mathbf{C}\in\mathbb{R}^{N\times 1}$ are projection matrices.
-To process discrete sequences, Mamba discretizes the continuous parameters $\mathbf{A}$ and $\mathbf{B}$ using a time scale parameter $\Delta$ and the zero-order hold (ZOH) principle, resulting in discretized parameters $\mathbf{\overline{A}}$ and $\mathbf{\overline{B}}$:
-$\mathbf{\overline{A}}$ $=\exp\left(\mathbf{\Delta A}\right),$ (10)
-$\mathbf{\overline{B}}$ $=\left(\mathbf{\Delta A}\right)^{-1}\left(\exp\left(\mathbf{\Delta A}\right)-\mathbf{I}\right)\cdot\mathbf{\Delta B}.$
-The discrete state-space equation with a step size of $\Delta$ becomes:
-$h_{t}$ $=\mathbf{\overline{A}}h_{t-1}+\mathbf{\overline{B}}x_{t},$ (11)
-$y_{t}$ $=\mathbf{C}h_{t}.$
-By iteratively expanding the hidden state $h_{t-1}$, Mamba derives a global convolution kernel $\mathbf{\overline{K}}\in\mathbb{R}^{L}$.
-This kernel is then used to compute the output $y$ through a convolution operation with the input $x$:
-$\mathbf{\overline{K}}$ $=\left(\mathbf{C\overline{B},C\overline{A}\overline{B},...,C\overline{A}^{\mathrm{L}-1}\overline{B}}\right),$ (12)
-$y$ $=x\otimes\mathbf{\overline{K}}.$
-Falcon Mamba 7B adopts a pure Mamba architecture, a departure from hybrid designs incorporating staggered attention.
-This deliberate choice aims to maintain the intrinsic linear scalability characteristic of Mamba models.
-To enhance adaptability, the model employs decoupled input embeddings and output weights.
-At its core, Falcon-Mamba features 64 layers of the Falcon-Mamba Mixer.
-Each Mixer layer integrates an SSM (State Space Model) module alongside in-projection and out-projection layers, RMS Norm, and a convolutional layer.
-Within the SSM module, the input is mapped to $\Delta$, $B$, and $C$ through a projection layer denoted as $x$-proj:
-$x\xrightarrow{x\text{-proj}}(\Delta,B,C)$
-where $x$ represents the input to the SSM module. Furthermore, another projection layer, $dt$-proj, discretizes $\Delta$:
-$\Delta\xrightarrow{dt\text{-proj}}\Delta_{discretized}$
-These discretized values—$\Delta_{discretized}$, $A$, $B$, $C$, and $D$—are then fed into the selective scanning module for processing.
-This architectural design of Falcon-Mamba fully enables the application of LoRA-Mixer specifically tailored for the projection layer.
-For a comprehensive understanding of the training process, please refer to.
-Appendix C Theoretical Justification of RSL Loss
-We provide a theoretical analysis of the proposed RSL loss and contrast it with the conventional auxiliary loss.
-Our goal is to demonstrate that RSL naturally promotes input-aware, expert-specialized routing with improved data efficiency.
-C.1 Preliminaries
-Let the router output a softmax distribution $p(x)=[p_{1}(x),\dots,p_{K}(x)]\in\Delta^{K-1}$ over $K$ experts for a token $x$, where
-$p_{i}(x)=\frac{\exp(G_{i}(x))}{\sum_{j=1}^{K}\exp(G_{j}(x))},\quad\sum_{i=1}^{K}p_{i}(x)=1.$ (13)
-We define the expected routing probability and top-1 selection frequency as:
-$\bar{p}_{i}$ $=\mathbb{E}_{x\sim\mathcal{D}}[p_{i}(x)],$ (14)
-$\bar{f}_{i}$ $=\mathbb{E}_{x\sim\mathcal{D}}\left[\mathbb{I}(i=\arg\max_{j}p_{j}(x))\right],$ (15)
-where $\bar{p}_{i}$ represents the average routing intention, and $\bar{f}_{i}$ represents the empirical usage under hard top-1 routing.
-C.2 Auxiliary Loss and Its Implicit Bias
-The standard auxiliary loss encourages load balancing by aligning the average routing with actual expert usage:
-$\mathcal{L}_{\text{aux}}=\alpha\sum_{i=1}^{K}\bar{p}_{i}\cdot\bar{f}_{i}.$ (16)
-Proposition 1 (Equilibrium of Auxiliary Loss).
-Under the constraint $\sum_{i=1}^{K}\bar{p}_{i}=1$, the minimum of $\mathcal{L}_{\text{aux}}$ is attained when $\bar{p}_{i}=\bar{f}_{i}$ for all $i$.
-Proof. Using the Lagrangian method:
-$\mathcal{L}=\sum_{i=1}^{K}\bar{p}_{i}\bar{f}_{i}-\lambda\left(\sum_{i=1}^{K}\bar{p}_{i}-1\right)$
-Taking partial derivatives:
-$\frac{\partial\mathcal{L}}{\partial\bar{p}_{i}}=\bar{f}_{i}-\lambda=0\Rightarrow\bar{f}_{i}=\lambda,\quad\forall i.$
-Thus, all $\bar{f}_{i}$ are equal, implying uniform distribution: $\bar{f}_{i}=\frac{1}{K}$, hence $\bar{p}_{i}=\frac{1}{K}$.
-This shows that the auxiliary loss alone biases the router toward uniform expert activation, regardless of input characteristics.
-C.3 RSL Loss: Entropy-Regularized Routing
-To promote more input-sensitive routing, we introduce an entropy-regularized objective:
-$\mathcal{L}_{\text{RSL}}=\mathcal{L}_{\text{aux}}-\lambda\cdot\mathbb{E}_{x\sim\mathcal{D}}[\mathcal{H}(p(x))],$ (17)
-where the token-level entropy is defined as:
-$\mathcal{H}(p(x))=-\sum_{i=1}^{K}p_{i}(x)\log p_{i}(x).$ (18)
-This term encourages the router to assign higher weights to fewer experts, effectively breaking uniformity and encouraging selective specialization.
-C.4 Gradient Analysis and Token-Awareness
-We derive the entropy gradient w.r.t. routing score $p_{i}(x)$:
-$\frac{\partial\mathcal{H}(p(x))}{\partial p_{i}(x)}$ $=-\log p_{i}(x)-1,\quad\text{subject to}\quad\sum_{i=1}^{K}p_{i}(x)=1,$ (19)
-$=-\log p_{i}(x)-1+\mu,$ (20)
-where $\mu$ is the Lagrange multiplier due to the simplex constraint.
-Thus, the total gradient of the RSL loss becomes:
-$\nabla_{p_{i}(x)}\mathcal{L}_{\text{RSL}}=\alpha\cdot\frac{\partial\bar{p}_{i}}{\partial p_{i}(x)}\cdot\bar{f}_{i}+\lambda(\log p_{i}(x)+1-\mu).$ (21)
-This shows that RSL introduces token-level gradient signals via $\log p_{i}(x)$, unlike the auxiliary loss, which propagates only global gradients.
-C.5 Token-Awareness via Routing Variance
-To quantify input-aware routing, we define:
-$\text{Var}_{x\sim\mathcal{D}}(p(x)):=\mathbb{E}_{x}\left[\|p(x)-\bar{p}\|^{2}\right].$ (22)
-We say the routing is token-aware if $\text{Var}(p(x))>\epsilon$ for some $\epsilon>0$.
-The auxiliary loss tends to reduce this variance (driving uniform routing), while RSL encourages high variance and peaked distributions aligned with input semantics.
-C.6 Conclusion
-The RSL loss incorporates an entropy-based regularizer that mitigates the uniformity bias of auxiliary loss.
-By injecting token-level gradient signals and promoting routing variance, RSL enables input-aware, specialized, and discriminative expert assignments.
-This property is especially beneficial in data-scarce regimes, where each token’s contribution to routing must be maximally leveraged.
-Note that RSL is fully compatible with auxiliary loss; it can be viewed as a strict generalization that stabilizes training while encouraging specialization.
+A.13 Balance loss visualization.
 Figure 6: Balance loss curve using RSL loss during training.
-Appendix D Balance loss visualization
 Figure 6 shows the changing trend of the Balance Loss when the RSL loss function is used during training.
@@
 Overall, this Balance Loss curve not only reflects the model’s ability to converge quickly, but also demonstrates the robustness of the training process, verifying the significant advantages of the RSL loss function in improving model performance and training efficiency.
-Appendix E Details of LoRA modules of Flan-T5.
+A.14 In-distribution experimental results.
+Here, we present experimental results for LoRA-Mixer and PHATGOOSE on five in-distribution experiments (GSM8K, ARC-E, ARC-C, CoLA, and SST2).
+The results are shown in Table 21.
+Both methods use the same LoRA adapter (Q/K/V/O) on LLaMA3-8B and are trained using the same data; their routing strategies differ: LoRA-Mixer uses RSL, while PHATGOOSE uses sigmoid gating based on cosine similarity.
+It can be seen that LoRA-Mixer outperforms PHATGOOSE across the board.
+Table 21: Comparison of Base, Phatgoose, and LoRA-Mixer across five datasets.
+Dataset Base Phatgoose LoRA-Mixer
+SST2 93.12 91.35 94.83
+CoLA 79.14 78.27 81.17
+GSM8K 57.92 58.87 62.66
+ARC-E 88.45 87.13 89.47
+ARC-C 78.65 75.94 79.89
+In order to verify the enhanced performances of individual expert after LoRA-Mixer optimization.
+We selected four tasks, GSM8K, SST2, CoLA and HumanEval, for experiments.
+The results are shown in Table LABEL:tab:expert_ablation_gain_with_lora_reordered.
+We can find that LoRAs optimized by LoRA-mixer exhibit improved performance, especially in the GSM8K task, where the performance improved by 5.36% after adding mathematical experts.
+This result confirms the enhanced ability of each individual LoRA expert after LoRA-Mixer optimization.
+A.15 Details of LoRA modules of Flan-T5.
+To better understand the Flan-T5 LoRAs experiment from the Internet, we introduce the corresponding LoRAs configuration information here, details as 22:
+Table 22: LoRA configuration details used in our experiments.
 Parameter Value
@@
 task_type SEQ_2_SEQ_LM
-Table 12: LoRA configuration details used in our experiments.
-Appendix F Experimental details
-Our experiments are conducted on a Linux workstation equipped with a single NVIDIA A800 80GB GPU and a 32-core Intel Xeon CPU.
-We use the AdamW optimizer with a learning rate of $1\times 10^{-5}$.
-For Transformer-based models, LoRA-Mixer is applied exclusively to the attention modules.
-For SSM-based models, LoRA-Mixer is integrated into the in, out, dt, and x projection layers.
-NeurIPS Paper Checklist
-  1. 1.
-    Claims
-  2. Question: Do the main claims made in the abstract and introduction accurately reflect the paper’s contributions and scope?
-  3. Answer: [Yes]
-  4.
-Justification: The abstract and introduction of the paper accurately reflect the contributions of the paper, such as proposing LoRA-Mixer, a highly flexible MoE framework, and verifying its effectiveness through a large number of experiments.
-  5. Guidelines:
-    • The answer NA means that the abstract and introduction do not include the claims made in the paper.
-    • The abstract and/or introduction should clearly state the claims made, including the contributions made in the paper and important assumptions and limitations.
-A No or NA answer to this question will not be perceived well by the reviewers.
-    • The claims made should match theoretical and experimental results, and reflect how much the results can be expected to generalize to other settings.
-    • It is fine to include aspirational goals as motivation as long as it is clear that these goals are not attained by the paper.
-  6. 2.
-    Limitations
-  7. Question: Does the paper discuss the limitations of the work performed by the authors?
-  8. Answer: [Yes]
-  9.
-Justification: Due to the length limitations of the article, we did not include a separate section in the main text to discuss the limitations of the article.
-We briefly discuss the limitations in the Appendix.
-  10. Guidelines:
-    • The answer NA means that the paper has no limitation while the answer No means that the paper has limitations, but those are not discussed in the paper.
-    • The authors are encouraged to create a separate "Limitations" section in their paper.
-    • The paper should point out any strong assumptions and how robust the results are to violations of these assumptions (e.g., independence assumptions, noiseless settings, model well-specification, asymptotic approximations only holding locally).
-The authors should reflect on how these assumptions might be violated in practice and what the implications would be.
-    • The authors should reflect on the scope of the claims made, e.g., if the approach was only tested on a few datasets or with a few runs.
-In general, empirical results often depend on implicit assumptions, which should be articulated.
-    • The authors should reflect on the factors that influence the performance of the approach.
-For example, a facial recognition algorithm may perform poorly when image resolution is low or images are taken in low lighting.
-Or a speech-to-text system might not be used reliably to provide closed captions for online lectures because it fails to handle technical jargon.
-    • The authors should discuss the computational efficiency of the proposed algorithms and how they scale with dataset size.
-    • If applicable, the authors should discuss possible limitations of their approach to address problems of privacy and fairness.
-    • While the authors might fear that complete honesty about limitations might be used by reviewers as grounds for rejection, a worse outcome might be that reviewers discover limitations that aren’t acknowledged in the paper.
-The authors should use their best judgment and recognize that individual actions in favor of transparency play an important role in developing norms that preserve the integrity of the community.
-Reviewers will be specifically instructed to not penalize honesty concerning limitations.
-  11. 3.
-    Theory assumptions and proofs
-  12. Question: For each theoretical result, does the paper provide the full set of assumptions and a complete (and correct) proof?
-  13. Answer: [Yes]
-  14. Justification: In the appendix, we provide a thorough theoretical analysis on why the RSL function is more effective than traditional auxiliary loss.
-  15. Guidelines:
-    • The answer NA means that the paper does not include theoretical results.
-    • All the theorems, formulas, and proofs in the paper should be numbered and cross-referenced.
-    • All assumptions should be clearly stated or referenced in the statement of any theorems.
-    • The proofs can either appear in the main paper or the supplemental material, but if they appear in the supplemental material, the authors are encouraged to provide a short proof sketch to provide intuition.
-    • Inversely, any informal proof provided in the core of the paper should be complemented by formal proofs provided in appendix or supplemental material.
-    • Theorems and Lemmas that the proof relies upon should be properly referenced.
-  16. 4.
-    Experimental result reproducibility
-  17.
-Question: Does the paper fully disclose all the information needed to reproduce the main experimental results of the paper to the extent that it affects the main claims and/or conclusions of the paper (regardless of whether the code and data are provided or not)?
-  18. Answer: [Yes]
-  19.
-Justification: The experimental section of our article fully describes the experimental configuration we used, such as the configuration of LoRA, the details of the dataset, the evaluation criteria used, etc.
-To ensure the reproducibility of the article, we will open source the code.
-  20. Guidelines:
-    • The answer NA means that the paper does not include experiments.
-    • If the paper includes experiments, a No answer to this question will not be perceived well by the reviewers: Making the paper reproducible is important, regardless of whether the code and data are provided or not.
-    • If the contribution is a dataset and/or model, the authors should describe the steps taken to make their results reproducible or verifiable.
-    • Depending on the contribution, reproducibility can be accomplished in various ways.
-For example, if the contribution is a novel architecture, describing the architecture fully might suffice, or if the contribution is a specific model and empirical evaluation, it may be necessary to either make it possible for others to replicate the model with the same dataset, or provide access to the model.
-In general.
-releasing code and data is often one good way to accomplish this, but reproducibility can also be provided via detailed instructions for how to replicate the results, access to a hosted model (e.g., in the case of a large language model), releasing of a model checkpoint, or other means that are appropriate to the research performed.
-    • While NeurIPS does not require releasing code, the conference does require all submissions to provide some reasonable avenue for reproducibility, which may depend on the nature of the contribution.
-For example
-      1. (a)
-        If the contribution is primarily a new algorithm, the paper should make it clear how to reproduce that algorithm.
-      2. (b)
-        If the contribution is primarily a new model architecture, the paper should describe the architecture clearly and fully.
-      3. (c)
-        If the contribution is a new model (e.g., a large language model), then there should either be a way to access this model for reproducing the results or a way to reproduce the model (e.g., with an open-source dataset or instructions for how to construct the dataset).
-      4. (d)
-        We recognize that reproducibility may be tricky in some cases, in which case authors are welcome to describe the particular way they provide for reproducibility.
-In the case of closed-source models, it may be that access to the model is limited in some way (e.g., to registered users), but it should be possible for other researchers to have some path to reproducing or verifying the results.
-  21. 5.
-    Open access to data and code
-  22. Question: Does the paper provide open access to the data and code, with sufficient instructions to faithfully reproduce the main experimental results, as described in supplemental material?
-  23. Answer: [Yes]
-  24.
-Justification: We will open source the code and submit it in the supplementary materials.
-As for the datasets, we have clearly stated them in the article.
-These datasets are open source and can fully guarantee the authenticity of the content of the article.
-  25. Guidelines:
-    • The answer NA means that paper does not include experiments requiring code.
-    • Please see the NeurIPS code and data submission guidelines (https://nips.cc/public/guides/CodeSubmissionPolicy) for more details.
-    • While we encourage the release of code and data, we understand that this might not be possible, so “No” is an acceptable answer.
-Papers cannot be rejected simply for not including code, unless this is central to the contribution (e.g., for a new open-source benchmark).
-    • The instructions should contain the exact command and environment needed to run to reproduce the results.
-See the NeurIPS code and data submission guidelines (https://nips.cc/public/guides/CodeSubmissionPolicy) for more details.
-    • The authors should provide instructions on data access and preparation, including how to access the raw data, preprocessed data, intermediate data, and generated data, etc.
-    • The authors should provide scripts to reproduce all experimental results for the new proposed method and baselines.
-If only a subset of experiments are reproducible, they should state which ones are omitted from the script and why.
-    • At submission time, to preserve anonymity, the authors should release anonymized versions (if applicable).
-    • Providing as much information as possible in supplemental material (appended to the paper) is recommended, but including URLs to data and code is permitted.
-  26. 6.
-    Experimental setting/details
-  27. Question: Does the paper specify all the training and test details (e.g., data splits, hyperparameters, how they were chosen, type of optimizer, etc.) necessary to understand the results?
-  28. Answer: [No]
-  29. Justification: Due to space limitations, we did not explore these details in detail in this paper.
-  30. Guidelines:
-    • The answer NA means that the paper does not include experiments.
-    • The experimental setting should be presented in the core of the paper to a level of detail that is necessary to appreciate the results and make sense of them.
-    • The full details can be provided either with the code, in appendix, or as supplemental material.
-  31. 7.
-    Experiment statistical significance
-  32. Question: Does the paper report error bars suitably and correctly defined or other appropriate information about the statistical significance of the experiments?
-  33. Answer: [Yes]
-  34. Justification: In order to reduce the error, the experimental results of our article are the average results obtained by repeating the experiment more than three times.
-  35. Guidelines:
-    • The answer NA means that the paper does not include experiments.
-    • The authors should answer "Yes" if the results are accompanied by error bars, confidence intervals, or statistical significance tests, at least for the experiments that support the main claims of the paper.
-    • The factors of variability that the error bars are capturing should be clearly stated (for example, train/test split, initialization, random drawing of some parameter, or overall run with given experimental conditions).
-    • The method for calculating the error bars should be explained (closed form formula, call to a library function, bootstrap, etc.)
-    • The assumptions made should be given (e.g., Normally distributed errors).
-    • It should be clear whether the error bar is the standard deviation or the standard error of the mean.
-    • It is OK to report 1-sigma error bars, but one should state it.
-The authors should preferably report a 2-sigma error bar than state that they have a 96% CI, if the hypothesis of Normality of errors is not verified.
-    • For asymmetric distributions, the authors should be careful not to show in tables or figures symmetric error bars that would yield results that are out of range (e.g. negative error rates).
-    • If error bars are reported in tables or plots, The authors should explain in the text how they were calculated and reference the corresponding figures or tables in the text.
-  36. 8.
-    Experiments compute resources
-  37. Question: For each experiment, does the paper provide sufficient information on the computer resources (type of compute workers, memory, time of execution) needed to reproduce the experiments?
-  38. Answer: [No]
-  39. Justification: Due to the length of the article, we do not discuss the computing resources in detail. We can briefly explain here that we use an A800 80G GPU for training and inference.
-  40. Guidelines:
-    • The answer NA means that the paper does not include experiments.
-    • The paper should indicate the type of compute workers CPU or GPU, internal cluster, or cloud provider, including relevant memory and storage.
-    • The paper should provide the amount of compute required for each of the individual experimental runs as well as estimate the total compute.
-    • The paper should disclose whether the full research project required more compute than the experiments reported in the paper (e.g., preliminary or failed experiments that didn’t make it into the paper).
-  41. 9.
-    Code of ethics
-  42. Question: Does the research conducted in the paper conform, in every respect, with the NeurIPS Code of Ethics https://neurips.cc/public/EthicsGuidelines?
-  43. Answer: [Yes]
-  44. Justification: Our paper complies with the ethical standards of the conference in all aspects and will not cause any adverse impact on society.
-  45. Guidelines:
-    • The answer NA means that the authors have not reviewed the NeurIPS Code of Ethics.
-    • If the authors answer No, they should explain the special circumstances that require a deviation from the Code of Ethics.
-    • The authors should make sure to preserve anonymity (e.g., if there is a special consideration due to laws or regulations in their jurisdiction).
-  46. 10.
-    Broader impacts
-  47. Question: Does the paper discuss both potential positive societal impacts and negative societal impacts of the work performed?
-  48. Answer: [No]
-  49. Justification: Due to the length limitation of the article, we did not discuss this aspect, but our research will not cause any adverse impact on society.
-  50. Guidelines:
-    • The answer NA means that there is no societal impact of the work performed.
-    • If the authors answer NA or No, they should explain why their work has no societal impact or why the paper does not address societal impact.
-    • Examples of negative societal impacts include potential malicious or unintended uses (e.g., disinformation, generating fake profiles, surveillance), fairness considerations (e.g., deployment of technologies that could make decisions that unfairly impact specific groups), privacy considerations, and security considerations.
-    • The conference expects that many papers will be foundational research and not tied to particular applications, let alone deployments.
-However, if there is a direct path to any negative applications, the authors should point it out.
-For example, it is legitimate to point out that an improvement in the quality of generative models could be used to generate deepfakes for disinformation.
-On the other hand, it is not needed to point out that a generic algorithm for optimizing neural networks could enable people to train models that generate Deepfakes faster.
-    • The authors should consider possible harms that could arise when the technology is being used as intended and functioning correctly, harms that could arise when the technology is being used as intended but gives incorrect results, and harms following from (intentional or unintentional) misuse of the technology.
-    • If there are negative societal impacts, the authors could also discuss possible mitigation strategies (e.g., gated release of models, providing defenses in addition to attacks, mechanisms for monitoring misuse, mechanisms to monitor how a system learns from feedback over time, improving the efficiency and accessibility of ML).
-  51. 11.
-    Safeguards
-  52.
-Question: Does the paper describe safeguards that have been put in place for responsible release of data or models that have a high risk for misuse (e.g., pretrained language models, image generators, or scraped datasets)?
-  53. Answer: [No]
-  54. Justification: The datasets and models we use are open source and safe. There is no risk.
-  55. Guidelines:
-    • The answer NA means that the paper poses no such risks.
-    • Released models that have a high risk for misuse or dual-use should be released with necessary safeguards to allow for controlled use of the model, for example by requiring that users adhere to usage guidelines or restrictions to access the model or implementing safety filters.
-    • Datasets that have been scraped from the Internet could pose safety risks. The authors should describe how they avoided releasing unsafe images.
-    • We recognize that providing effective safeguards is challenging, and many papers do not require this, but we encourage authors to take this into account and make a best faith effort.
-  56. 12.
-    Licenses for existing assets
-  57.
-Question: Are the creators or original owners of assets (e.g., code, data, models), used in the paper, properly credited and are the license and terms of use explicitly mentioned and properly respected?
-  58. Answer: [Yes]
-  59. Justification: The models and datasets we use are open source and have obtained asset licenses from their authors.
-  60. Guidelines:
-    • The answer NA means that the paper does not use existing assets.
-    • The authors should cite the original paper that produced the code package or dataset.
-    • The authors should state which version of the asset is used and, if possible, include a URL.
-    • The name of the license (e.g., CC-BY 4.0) should be included for each asset.
-    • For scraped data from a particular source (e.g., website), the copyright and terms of service of that source should be provided.
-    • If assets are released, the license, copyright information, and terms of use in the package should be provided.
-For popular datasets, paperswithcode.com/datasets has curated licenses for some datasets.
-Their licensing guide can help determine the license of a dataset.
-    • For existing datasets that are re-packaged, both the original license and the license of the derived asset (if it has changed) should be provided.
-    • If this information is not available online, the authors are encouraged to reach out to the asset’s creators.
-  61. 13.
-    New assets
-  62. Question: Are new assets introduced in the paper well documented and is the documentation provided alongside the assets?
-  63. Answer: [No]
-  64. Justification: We have not provided any explanation for this aspect and our model will also be open source.
-  65. Guidelines:
-    • The answer NA means that the paper does not release new assets.
-    • Researchers should communicate the details of the dataset/code/model as part of their submissions via structured templates. This includes details about training, license, limitations, etc.
-    • The paper should discuss whether and how consent was obtained from people whose asset is used.
-    • At submission time, remember to anonymize your assets (if applicable). You can either create an anonymized URL or include an anonymized zip file.
-  66. 14.
-    Crowdsourcing and research with human subjects
-  67.
-Question: For crowdsourcing experiments and research with human subjects, does the paper include the full text of instructions given to participants and screenshots, if applicable, as well as details about compensation (if any)?
-  68. Answer: [No]
-  69. Justification: Our experiments do not contain any experimental data related to humans and therefore need not be discussed.
-  70. Guidelines:
-    • The answer NA means that the paper does not involve crowdsourcing nor research with human subjects.
-    • Including this information in the supplemental material is fine, but if the main contribution of the paper involves human subjects, then as much detail as possible should be included in the main paper.
-    • According to the NeurIPS Code of Ethics, workers involved in data collection, curation, or other labor should be paid at least the minimum wage in the country of the data collector.
-  71. 15.
-    Institutional review board (IRB) approvals or equivalent for research with human subjects
-  72.
-Question: Does the paper describe potential risks incurred by study participants, whether such risks were disclosed to the subjects, and whether Institutional Review Board (IRB) approvals (or an equivalent approval/review based on the requirements of your country or institution) were obtained?
-  73. Answer: [No]
-  74. Justification: Our experiments did not involve any experimental subjects.
-  75. Guidelines:
-    • The answer NA means that the paper does not involve crowdsourcing nor research with human subjects.
-    • Depending on the country in which research is conducted, IRB approval (or equivalent) may be required for any human subjects research.
-If you obtained IRB approval, you should clearly state this in the paper.
-    • We recognize that the procedures for this may vary significantly between institutions and locations, and we expect authors to adhere to the NeurIPS Code of Ethics and the guidelines for their institution.
-    • For initial submissions, do not include any information that would break anonymity (if applicable), such as the institution conducting the review.
-  76. 16.
-    Declaration of LLM usage
-  77.
-Question: Does the paper describe the usage of LLMs if it is an important, original, or non-standard component of the core methods in this research?
-Note that if the LLM is used only for writing, editing, or formatting purposes and does not impact the core methodology, scientific rigorousness, or originality of the research, declaration is not required.
-  78. Answer: [Yes]
-  79. Justification: The LLMs we use are all open source and We use the large language model for grammar polishing.
-  80. Guidelines:
-    • The answer NA means that the core method development in this research does not involve LLMs as any important, original, or non-standard components.
-    • Please refer to our LLM policy (https://neurips.cc/Conferences/2025/LLM) for what should or should not be described.
+A.16 Explanation of RSL’s suboptimal performance with 4k data.
+RSL’s temporary performance fluctuations on 4k datasets can be explained by its design philosophy.
+RSL integrates an entropy regularizer to balance expert specialization and load balancing.
+On small datasets (1k-3k instances), this mechanism helps the router quickly focus on the most relevant experts by avoiding overly uniform activation functions, thus outperforming common auxiliary loss functions that tend to enforce an indiscriminate balance.
+When the number of instances reaches 4k, RSL begins to explore finer-grained expert tasks (consistent with its goal of input-aware routing).
+However, this phase can lead to temporary instability, as the data volume is sufficient to trigger exploration but not yet rich enough to fully stabilize finer routing patterns.
+In contrast, the original auxiliary loss function maintains a stable but suboptimal uniform activation function, showing a temporary advantage.
+However, as data size increases further, RSL’s advantage increases.
+This is because the auxiliary loss becomes overly averaged and loses input awareness, resulting in less accurate expert selection and suboptimal performance.
+A.17 EXCESSIVE AVERAGING OF AUXILIARY LOSSES.
+The standard auxiliary loss in mixture-of-experts is designed to encourage load balancing by aligning the average routing probability $\bar{p}_{i}$ with the average usage frequency $\bar{f}_{i}$:
+$L_{\text{aux}}=\alpha\sum_{i=1}^{K}\bar{p}_{i}\,\bar{f}_{i},\qquad\sum_{i=1}^{K}\bar{p}_{i}=1.$ (20)
+Here $\bar{p}_{i}=\mathbb{E}[p_{i}(x)]$ and $\bar{f}_{i}=\mathbb{E}[\mathbf{1}\{i=\arg\max_{j}p_{j}(x)\}]$.
+Since $\bar{f}_{i}$ depends on hard top-1 routing and is not differentiable, implementations usually adopt a smooth surrogate $\bar{s}_{i}$, e.g.
+softmax usage with temperature $\tau>0$.
+With this surrogate, the alignment loss is often combined or replaced with a quadratic balancing form:
+$L_{\text{bal}}=\alpha\sum_{i=1}^{K}\bar{p}_{i}^{2},\qquad\sum_{i=1}^{K}\bar{p}_{i}=1,$ (21)
+which penalizes deviations from equal usage. Consider the Lagrangian
+$\mathcal{L}(\bar{p},\lambda)=\sum_{i=1}^{K}\bar{p}_{i}^{2}-\lambda\Big(\sum_{i=1}^{K}\bar{p}_{i}-1\Big).$ (22)
+Setting $\partial\mathcal{L}/\partial\bar{p}_{i}=0$ yields $2\bar{p}_{i}-\lambda=0\;\Rightarrow\;\bar{p}_{i}=\lambda/2$ for all $i$. Summing over $i$ gives $\lambda=2/K$, hence
+$\bar{p}_{i}^{\star}=\tfrac{1}{K},\qquad\forall i.$ (23)
+The unique minimizer of $L_{\text{bal}}$ is the uniform distribution, showing that traditional auxiliary losses inherently bias the router towards equal activation and thus hinder expert specialization.
+A.18 Use of LLMs.
+We use LLMs as a basemodel for our experiments, and we also use LLMs for syntax polishing and checking.